### v2

# An explicit local geometric Langlands for supercuspidal representations: the toral case

Abstract.
We formulate a conjecture on local geometric Langlands for supercuspidal representations using Yu’s data and Feigin-Frenkel isomorphism. We refine our conjecture for a large family of regular supercuspidal representations defined by Kaletha, and then confirm the conjecture for toral supercuspidal representations of Adler whose Langlands parameters turn out to be exactly all the irreducible isoclinic connections. As an application, we establish the conjectural correspondence between global Airy connections for reductive groups and the family of Hecke eigensheaves constructed by Jakob-Kamgarpour-Yi.
1. Introduction
1.1. Supercuspidal representations and Yu data
In the seminal work of Jiu-Kang Yu [Yu], he proposed a construction of supercuspidal representations of tame $p$-adic groups, which turns out to be exhaustive [Kim, FintzenType] under some assumptions on the characteristic of the residue field. Let $G$ be a connected reductive group over a non-archimedean local field $F$ with characteristic $p\neq 2$. Assume $G$ splits over a tamely ramified extension of $F$. The input of the construction is a Yu datum consisting of a tuple
$((G_{i})_{1\leq i\leq n},x,(r_{i})_{1\leq i\leq n},\rho,(\phi_{i})_{1\leq i\leq n}),$
where $G=G_{0}\supset G_{1}\supset\cdots\supset G_{n}$ are twisted Levi subgroups that split over a tamely ramified extension of $F$, $x$ is a point in the Bruhat-Tits building of $G_{n}$, $r_{1}>r_{2}>\cdots>r_{n}>0$ are real numbers, $\rho$ is an irreducible representation of the reductive quotient of the parahoric subgroup of $G_{n}$ defined by $x$, and $\phi_{i}$ is a character of $G_{i}(F)$ of depth $r_{i}$, where these data need to satisfy some extra assumptions. From such a datum, Yu constructed a compact open subgroup $J\subset G(F)$ and a representation $\tilde{\phi}$ such that the compact induced representation
(1) $\mathrm{c-ind}^{G(F)}_{J}\tilde{\phi}$
is irreducible, thus supercuspidal [Yu, FintzenTame].
On the other hand, in a series of works of Kaletha [KalehtaSimple, KalethaEpipelagic, KalethaRegular, KalethaPacket], he constructed L-packets for supercuspidal L-parameters, providing a potential explicit local Langlands correspondence for supercuspidal representations. In the case of epipelagic representations, he proved that for $G=\mathrm{GL}_{n}$, his bijection coincides with the local Langlands correspondence of Bushnell-Henniart, see [KalethaEpipelagic, §6]. ver, to the knowledge of the author, beyond the above situation, there is no known compatibility between Kaletha’s construction with any local Langlands correspondence that has been established.
1.2. A conjectural supercuspidal local geometric Langlands
The goal of this article is to propose a compatibility between Kaletha’s construction with local geometric Langlands of Frenkel-Gaitsgory [FGLocal] in the de Rham setting, which we confirm for toral supercuspidal representations.
1.2.1. Local geometric Langlands
Let $F=\mathbb{C}(\!(t)\!)$ and $\mathcal{O}=\mathbb{C}[\![t]\!]$, then $D^{\times}=\mathrm{Spec}F$ and $D=\mathrm{Spec}\mathcal{O}$ are punctured formal disk and formal disk. Let $G/\mathbb{C}$ be a simply-connected simple algebraic group with Lie algebra $\mathfrak{g}$, and let $\check{G}$, $\check{\mathfrak{g}}$ be its dual group and dual Lie algebra. In the de Rham setting, a local L-parameter is a formal $\check{G}$-connection $\nabla$ over $D^{\times}$. It is proved by Frenkel-Zhu [FZOper] that any formal $\check{G}$-connection $\nabla$ can be equipped with a so-called oper structure. Roughly speaking, a $\check{G}$-oper consists of a $\check{G}$-bundle $\mathcal{F}$, a flat connection $\nabla$ on $\mathcal{F}$, and a Borel reduction $\mathcal{F}_{B}$ of $\mathcal{F}$ satisfying a strong transversality condition, see [BD] for the precise definition. Denote by $\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})$ the space of $\check{G}$-opers on $D^{\times}$ and by $\mathrm{Loc}_{\check{G}}(D^{\times})$ the stack of $\check{G}$-connections on $D^{\times}$. We have surjection
(2) $p:\ \mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})\rightarrow\mathrm{Loc}_{\check{G}}(D^{\times}).$
Let $\widehat{\mathfrak{g}}$ be the affine Kac-Moody algebra for $\mathfrak{g}$ at critical level. It is a particular central extension of the loop algebra $\mathfrak{g}(\!(t)\!)$ by central element $\mathbf{1}$. Let $\widehat{U}(\mathfrak{g})$ be the completed universal enveloping algebra of $\widehat{\mathfrak{g}}$ in which the central element $\mathbf{1}$ is identified with the unit $1$. Its modules are smooth representations of $\widehat{\mathfrak{g}}$ at critical level, which we denote by $\widehat{\mathfrak{g}}-\mathrm{mod}$. Let $\mathfrak{Z}$ be the center of $\widehat{U}(\mathfrak{g})$. The Feigin-Frenkel isomorphism states that there is an isomorphism
(3) $\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})\simeq\mathrm{Spec}\mathfrak{Z}$
compatible with a $\mathbb{G}_{m}$-action and certain Poisson structures on both sides.
Given a $\check{G}$-connection $\nabla$ on $D^{\times}$, we take any oper structure $\chi\in p^{-1}(\nabla)$ above it. Via (3), it defines a character $\chi:\mathfrak{Z}\rightarrow\mathbb{C}$. Let $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ be the category of smooth $\widehat{\mathfrak{g}}$-modules on which $\mathfrak{Z}$ acts via $\chi$. This category is what Frenkel and Gaitsgory proposed to be the local Langlands correspondence of $\nabla$. More precisely, the loop group $LG=G(\!(t)\!)$ naturally acts on $\widehat{\mathfrak{g}}$ [FrenkelBook, page 29], inducing an action of $LG$ on the $\widehat{\mathfrak{g}}$-modules by pre-composing the action on $\widehat{\mathfrak{g}}$. This action commutes with the action of center $\mathfrak{Z}$, so that $LG$ acts on category $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$, and moreover on its Grothendieck group $K_{0}(\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi})$. The resulting representation of $LG$ is the analog of the admissible representation in the classical local Langlands correspondence for $p$-adic groups.
It is conjectured in [FGLocal] that for different oper structures $\chi,\chi^{\prime}$ over the same connection $\nabla$, the categories $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$, $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi^{\prime}}$ are equivalent. Some corollaries of this conjecture have been confirmed in the spherical and tamely ramified cases [FGSpherical, FGFlag].
1.2.2. An explicit correspondence for supercuspidal representations
Replacing $p$-adic group with loop group in a Yu datum, we obtain a $K$-type $(J,\tilde{\phi})$ where $J\subset LG$ and $\tilde{\phi}$ is a representation of $J$ [^1] The use of Weil representations in the construction of $\tilde{\phi}$ needs to be replaced by extending $J$ by some Lagrangian subspaces.. Choosing the point $x$ in the closure of the fundamental alcove, we can let $J$ be contained in the positive loop group $L^{+}G=G[\![t]\!]$. Denote by $\mathfrak{j}$ the Lie algebra of $J$, and still denote by $\tilde{\phi}$ the associated representation of $\mathfrak{j}$. Since $\mathfrak{j}\subset\mathfrak{g}[\![t]\!]$, $\mathfrak{j}+\mathbb{C}\mathbf{1}\subset\widehat{\mathfrak{g}}$ is a subalgebra. Consider induced $\widehat{\mathfrak{g}}$-representation
(4) $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}:=\mathrm{Ind}_{\mathfrak{j}+\mathbb{C}\mathbf{1}}^{\widehat{\mathfrak{g}}}\tilde{\phi}.$
Denote its central support by
(5) $\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}:=\mathrm{Im}(\mathfrak{Z}\rightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}),\qquad\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi}):=\mathrm{Spec}\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}\hookrightarrow\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times}).$
Conjecture 1.
  • (i)
    There exists a unique irreducible formal $\check{G}$-connection $\nabla_{\tilde{\phi}}$ such that $\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})=p^{-1}(\nabla_{\tilde{\phi}})$, i.e. all the oper structures on $\nabla_{\tilde{\phi}}$.
  • (ii)
    For different Yu data that induce isomorphic supercuspidal representations of the $p$-adic group, the corresponding $K$-types $(J,\tilde{\phi})$, $(J^{\prime},\tilde{\phi}^{\prime})$ for loop group induce the same central support: $\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})=\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j}^{\prime},\tilde{\phi}^{\prime})$. In particular, their corresponding formal $\check{G}$-connections in (i) are isomorphic: $\nabla_{\tilde{\phi}}\simeq\nabla_{\tilde{\phi}^{\prime}}$.
  • (iii)
    Every irreducible formal $\check{G}$-connection arises as some $\nabla_{\tilde{\phi}}$.
  • (iv)
    The above correspondence between $\nabla_{\tilde{\phi}}$ and Yu data coincides with the supercuspidal L-packets constructed by Kaletha [KalethaPacket].
Remark 2.
  • (i)
    The conditions for different Yu data to give isomorphic supercuspidal representations are given by Hakim and Murnaghan [HM].
  • (ii)
    For epipelagic representations, a special family of supercuspidal representations of depth $\frac{1}{m}$ for $m$ a regular elliptic number, part (i) of the conjecture has been essentially proved in [CYTheta, §4, §5].
We will refine the above conjecture by introducing the notion of refined leading terms for both formal connections and $K$-types $(J,\tilde{\phi})$ in §2 and §3. In §4, we use refined leading terms to give a refinement of Conjecture 1 for a large family of regular supercuspidal representations. These supercuspidal representations are defined and studied by Kaletha [KalethaRegular]. We will see that the refinement of the conjecture leads to a potential approach to its proof, which we plan to revisit in the future works.
Our formulation of local geometric Langlands for the above representations is explicit in the following sense: the Langlands parameter $\nabla_{\tilde{\phi}}$ is conjectured to be the gauge equivalence class of a ‘minimal’ oper canonical form, where the coefficients of the equation are directly determined by the character $\tilde{\phi}$ as in Remark 16.
1.2.3. Evidence to Frenkel-Gaitsgory conjecture
Conjecture 1 provides evidence to Frenkel-Gaitsgory’s conjecture that $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ is independent of the choice of the oper structure. Let $\chi$ be an arbitrary oper structure on $\nabla_{\tilde{\phi}}$. By the definition of $\nabla_{\tilde{\phi}}$, $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}/\ker\chi$ is a $(J,\tilde{\phi})$-equivariant object in the category $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$. Suppose $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}/\ker\chi$ is nonzero. Then it provides a nonzero $(J,\tilde{\phi})$-eigenvector in the $LG$ representation $\pi_{\chi}=K_{0}(\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi})$ on the Grothendieck group. Heuristically, if we pretend as if $LG$ behaves the same as a $p$-adic group and formally consider the induced representation $\mathrm{c-ind}^{G(F)}_{J}\tilde{\phi}$, it is supposed to be irreducible, therefore supercuspidal. We obtain a nonzero map
(6) $\iota:\mathrm{c-ind}^{G(F)}_{J}\tilde{\phi}\rightarrow\pi_{\chi}.$
Since $\mathrm{c-ind}^{G(F)}_{J}\tilde{\phi}$ is irreducible, $\iota$ is isomorphic if $\pi_{\chi}$ is irreducible. In particular, $\pi_{\chi}$ depends only on $(J,\tilde{\phi})$ and is independent of the choice of oper structure $\chi$ on $\nabla_{\tilde{\phi}}$.
1.3. Isoclinic formal connections and toral supercuspidal representations
In this article, we confirm Conjecture 1.(i)-(iii) for toral supercuspidal representations, a family constructed by Adler [Adler] and named by Kaletha [KalethaRegular]. The Yu data of such supercuspidal representations contain only one proper twisted Levi subgroup given by an elliptic torus. This family of supercuspidal representations contain epipelagic representations and can attain arbitrary large depth.
It turns out that their L-parameters $\nabla_{\tilde{\phi}}$ defined via Conjecture 1 are exactly all the irreducible isoclinic connections, i.e. those irreducible formal $\check{G}$-connections whose canonical forms have regular semisimple leading terms, see §5 for details. Such formal connections have been studied in [JYDeligneSimpson], where the authors determine the conditions for the existence of a $\check{G}$-connection on $\mathbb{P}^{1}-\{0,\infty\}$ with one regular singularity at 00 with fixed monodromy class and one irregular singularity at $\infty$ with fixed isoclinic irregular type. See §5.1 for a review.
We summarize our main result as follows.
Theorem 3.
Conjecture 1.(i)-(iii) hold for toral supercuspidal representations. The L-parameters $\nabla_{\tilde{\phi}}$ are all the irreducible isoclinic foraml connections.
We will prove a refined statement in §7, see Lemma 35, Theorem 36, and Remark 37. An explicit equation of $\nabla_{\tilde{\phi}}$ in terms of $\tilde{\phi}$ will be given by (43), see Theorem 36.
1.4. Application to Airy connections
Let $h$ be the Coxeter number of $\check{G}$. Let $t$ be a coordinate at $0\in\mathbb{P}^{1}$. The isoclinic formal connections with slope $\frac{1+h}{h}$ over $D^{\times}$ can be globalized to $\check{G}$-connections over $\mathbb{P}^{1}-\{0\}$ called Airy $\check{G}$-connections, c.f. [HJRigid, KSRigid, JKY]. Such connections greatly generalize the classical Airy equation $y^{\prime\prime}=zy$ for $z=t^{-1}$. In [KSRigid], the authors constructed a family of Airy $\check{G}$-connections explicitly as follows.
Fix a root system of $\check{\mathfrak{g}}$ with simple roots. Let $p_{-1}$ be the sum of a set of basis of negative simple root subspaces, and $p_{n}$ a basis of the highest root subspace. Then
(7) $\nabla=\mathrm{d}+(t^{-2}p_{-1}+t^{-3}p_{n})\mathrm{d}t$
defines a family of Airy $\check{G}$-connections, see [KSRigid, equation (2)]. We give equations of arbitrary Airy $\check{G}$-connections in Proposition 38.
In Theorem 41 we show that the eigenvalues of the Hecke eigensheaves constructed in [JKY] are Airy $\check{G}$-connections, generalizing [JKY, Corollary 36] for the case of general linear groups. Our argument also provides a new proof of the physical rigidity of those Airy connections appearing as eigenvalues of these Hecke eigensheaves, partially recovering [HJRigid, Theorem 5.5.2].
1.5. Future directions
1.5.1. Applications in positive characteristic
The $K$-types $(J,\tilde{\phi})$ of toral supercuspidal representations of small depth have appeared as part of various rigid automorphic data. These data give rise to $\check{G}$-local systems on punctured projective line in both positive characteristic and characteristic zero setting. For $\ell$-adic sheaves on punctured projective line over a finite field, their $p$-adic companions can be realized as $F$-iscocrystals on a rigid space. In a joint work in progress with Daxin Xu, we study the $p$-adic local systems over Robba ring whose underlying connection is isoclinic. As an application, we combine the above with results in this article and [CYTheta] to establish some conjectural properties of the $\ell$-adic $\check{G}$-local systems constructed as eigenvalues of those Hecke eigensheaves in [KXY, YunEpipelagic]. In particular, we will deduce their $p$-adic physical rigidity and compute their Swan conductors.
1.5.2. Regular supercuspidal representations
Regular supercuspidal representations are generalization of toral supercuspidal representations that arise from a much simpler form of data than Yu data consisting of an elliptic maximal torus $S$ and certain character $\phi:S(F)\rightarrow\mathbb{C}^{\times}$. Kaletha associated Yu data to such pairs, which give rise to the regular supercuspidal representations. Also, he constructed L-packets for such representations. Our next step towards an explicit local geometric Langlands aims to prove Conjecture 1 for regular supercuspidal representations. A refined conjecture will be given in Conjecture 14.
1.5.3. Comparison with a wildly ramified Langlands
The results of this article should be closely related to the work of Bezrukavnikov-Boixeda-McBreen-Yun [BBMY]. In §5 of loc. cit., the authors proposed a conjectural wildly ramified global Langlands correspondece. At the unique wildly ramified place, the level structure they imposed is exactly $(J,\tilde{\phi})$ for a toral supercuspidal representation. We hope to investigate the application of our local results to their conjecture.
1.6. Organization of the article
In §2 and §3, we review the canonical forms of formal connections and Yu data of supercuspidal representations, and introduce their refined leading terms respectively. In §4, we refine our conjecture for a large family of regular supercuspidal representations. In §5, we determine oper structures on isoclinic connections. In §6, we review the construction of toral supercuspidal representations and compute the corresponding central supports. In §7, we prove the refined conjecture for irreducible isoclinic connections. In §8, we discuss applications on Airy $\check{G}$-connections.
Notations
Let $G$ be a simply-connected simple algebraic group over $\mathbb{C}$. Let $T\subset B\subset G$ be a maximal torus and Borel subgroup of $G$. Let $\mathfrak{t}\subset\mathfrak{b}\subset\mathfrak{g}$ be their Lie algebras. Let $\check{T}\subset\check{B}\subset\check{G}$ be the dual maximal torus and a fixed Borel in the dual group, and $\check{\mathfrak{t}}\subset\check{\mathfrak{b}}\subset\check{\mathfrak{g}}$ their Lie algebras. Let $n$ be the rank of $G$. Denote by $\Phi(G,T),\check{\Phi}(G,T),\Phi^{+}(G,T),\check{\Phi}^{+}(G,T),\Delta_{G},\check{\Delta}_{G}$ the (co)roots, positive (co)roots, simple (co)roots of $G$, and similarly for $\check{G}$ and Levi subgroups.
Denote by $\check{\rho}:=\check{\rho}_{\check{G}}$ the half sum of positive coroots of $\check{G}$. Let $\check{\mathfrak{g}}^{rs}$ be the locus of regular semisimple elements. For any subspace $\mathfrak{h}\subset\check{\mathfrak{g}}$, $\mathfrak{h}^{rs}:=\mathfrak{h}\cap\check{\mathfrak{g}}^{rs}$.
Let $W_{G}\simeq W_{\check{G}}$ be the Weyl group acting on $\check{\mathfrak{t}}$. We say an element $w\in W$ is regular if it has an eigenvector $v\in\check{\mathfrak{t}}$ that is not fixed by any element of $W$. The orders of regular elements are called regular numbers of $W$.
Let $F=\mathbb{C}(\!(t)\!)$, $\mathcal{O}=\mathbb{C}[\![t]\!]$. Let $\overline{F}$ be the algebraic closure of $F$. Let $D^{\times}=\mathrm{Spec}F\subset D=\mathrm{Spec}\mathcal{O}$ be punctured formal disk and formal disk.
Acknowledgement
The author thanks Konstantin Jakob and Daxin Xu for helpful discussions, thanks Zhiwei Yun for providing comments on a draft of the paper, and thanks Will Sawin for pointing out a mistake in §3.1.2. The research of Lingfei Yi is supported by the grant No. JIH1414033Y of Fudan University.
2. Refined leading terms of formal connections
2.1. Canonical forms
A formal $\check{G}$-connection $\nabla$ on $D^{\times}$ is a $\check{G}(\!(t)\!)$-gauge equivalence class of an equation of the following form:
$\mathrm{d}+(A_{-r}t^{-r}+A_{-r+1}t^{-r+1}+\cdots)\mathrm{d}t,\quad A_{i}\in\check{\mathfrak{g}},\ r\in\mathbb{Z},$
where the gauge transform of $g\in\check{G}(\!(t)\!)$ acts by
$g\cdot\nabla=\mathrm{d}+\mathrm{Ad}_{g}(A_{-r}t^{-r}+A_{-r+1}t^{-r+1}+\cdots)\mathrm{d}t-g^{-1}(\mathrm{d}g).$
The main theorem of [BV, §9.5] tells us that a formal connection $\nabla$ can always be $\check{G}(\overline{F})$-gauge transformed into the following form:
(8) $\mathrm{d}+(D_{1}t^{-r_{1}}+\cdots+D_{k}t^{-r_{k}}+D_{k+1})\frac{\mathrm{d}t}{t},$
where $r_{1}>r_{2}>\cdots>r_{k}>0$ are rational numbers, $r_{k+1}=0$, $D_{i}\in\check{\mathfrak{g}}$ are mutually commutative nonzero elements(if exists), and $D_{i}$’s are all semisimple for $i<0$. This is called a canonical form of $\nabla$. Any two canonical forms of $\nabla$ must have the same $r_{i}$’s, and the tuples $(D_{1},...,D_{k},\exp(2\pi\sqrt{-1}D_{k+1}))$ is unique up to conjugation by $\check{G}$. In particular, $\nu:=r_{1}$ is called the slope of $\nabla$. By [BV, §9.8 Lemma 1], we can choose the canonical form to be weakly $\mathbb{Z}$-reduced, i.e. the semisimple part $D_{k+1,s}$ of $D_{k+1}$ satisfies that for any $m\in\mathbb{Z}-\{0\}$, $D_{k+1,s}$ and $\exp(2\pi\sqrt{-1}mD_{k+1,s})$ have the same centralizer in the common centralizer $\check{\mathfrak{g}}_{D_{1},...,D_{k}}=\cap_{i=1}^{k}\mathfrak{z}_{\check{\mathfrak{g}}}(D_{i})$.
Conversely, given a weakly $\mathbb{Z}$-reduced form $\nabla$ over $\overline{F}$ as in (8), [BV, §9.8 Proposition] shows that such a form is the canonical form of a formal connection over $F$ if and only if there exists integer $b\geq 1$ and $\theta\in\check{G}$ satisfying $\theta^{b}=1$, $br_{j}\in\mathbb{Z}$, $\mathrm{Ad}_{\theta}(D_{i})=\exp(-2\pi\sqrt{-1}r_{i})D_{i}$ for all $i$.
2.2. Refined leading terms
2.2.1.
The nonzero semisimple coefficient $D_{1}\neq 0$ of (8) is called its leading term. In the following we define refined leading terms that encode more information from the connection.
We begin with a weakly $\mathbb{Z}$-reduced canonical form $\nabla$ as in (8). Fix once and for all a Cartan subalgebra $\check{\mathfrak{h}}\subset\check{\mathfrak{g}}_{D_{1},...,D_{k}}$ contained in the common centralizer of the irregular part. Let $\check{H}$ be the corresponding maximal torus in $\check{G}$. For $1\leq i\leq k$, let $\check{M}_{i}=C_{\check{G}}(D_{1},...,D_{i})$ be the common centralizer. We also let $\check{M}_{k+1}:=\check{H}$. Since $D_{i}$’s are mutually commuting semisimple elements of $\check{\mathfrak{h}}$, $\check{M}_{i}$ is the centralizer of the minimal subtorus whose Lie algebra contains $D_{1},...,D_{i}$, thus a connected Levi subgroup containing $\check{H}$. Denote its Lie algebra by $\check{\mathfrak{m}}_{i}$, the center of $\check{\mathfrak{m}}_{i}$ by $\check{\mathfrak{z}}_{i}$, the derived subalgebra by $\check{\mathfrak{m}}_{i}^{\mathrm{der}}=[\check{\mathfrak{m}}_{i},\check{\mathfrak{m}}_{i}]$. We take convention $\check{\mathfrak{m}}_{0}=\check{\mathfrak{g}},\check{\mathfrak{z}}_{0}=0$. The Levi decomposition gives $\check{\mathfrak{m}}_{i}=\check{\mathfrak{z}}_{i}\oplus\check{\mathfrak{m}}_{i}^{\mathrm{der}}$, where $\check{\mathfrak{z}}_{i}\subset\check{\mathfrak{h}}$. Note $D_{1},...,D_{i}\in\check{\mathfrak{z}}_{i}$, but $\check{\mathfrak{z}}_{i}$ may not be spanned by $D_{j}$’s, for instance when $X_{1}$ is regular semisimple. We have filtrations
$\check{\mathfrak{g}}=\check{\mathfrak{m}}_{0}\supset\check{\mathfrak{m}}_{1}\supset\check{\mathfrak{m}}_{2}\supset\cdots\supset\check{\mathfrak{m}}_{k}\supset\check{\mathfrak{m}}_{k+1}=\check{\mathfrak{h}},\qquad 0=\check{\mathfrak{z}}_{0}\subset\check{\mathfrak{z}}_{1}\subset\check{\mathfrak{z}}_{2}\subset\cdots\subset\check{\mathfrak{z}}_{k}\subset\check{\mathfrak{z}}_{k+1}=\check{\mathfrak{h}}.$
Denote $\check{\mathfrak{a}}_{i}:=\check{\mathfrak{z}}_{i}\cap\check{\mathfrak{m}}_{i-1}^{\mathrm{der}}$. Here $\check{\mathfrak{a}}_{1}=\check{\mathfrak{z}}_{1}$. Since $\check{\mathfrak{z}}_{i-1}\subset\check{\mathfrak{z}}_{i}$, the Levi decomposition $\check{\mathfrak{z}}_{i}\subset\check{\mathfrak{m}}_{i-1}=\check{\mathfrak{z}}_{i-1}\oplus\check{\mathfrak{m}}_{i-1}^{\mathrm{der}}$ induces decompositions
$\check{\mathfrak{z}}_{i}=\check{\mathfrak{z}}_{i-1}\oplus\check{\mathfrak{a}}_{i},\qquad\check{\mathfrak{z}}_{i}=\bigoplus_{1\leq j\leq i}\check{\mathfrak{a}}_{j}.$
Denote the projection of $D_{i}\in\check{\mathfrak{z}}_{i}$ to $\check{\mathfrak{a}}_{i}$ by $X_{i}$. Denote
$I=\{1\leq i\leq k\mid X_{i}\neq 0\}=\{i_{1}<i_{2}<\cdots<i_{p}\},\quad R=\{r_{i}\mid i\in I\},\quad X=(X_{i})_{i\in I}.$
Note that $i_{1}=1$. Equivalently, $I$ is the set of breaks of the Levi subalgebra filtration:
$\check{\mathfrak{m}}_{1}=\check{\mathfrak{m}}_{i_{1}}=\check{\mathfrak{m}}_{i_{1}+1}=\cdots=\check{\mathfrak{m}}_{i_{2}-1}\supsetneq\check{\mathfrak{m}}_{i_{2}}=\cdots=\check{\mathfrak{m}}_{i_{3}-1}\supsetneq\cdots\supsetneq\check{\mathfrak{m}}_{i_{p}}=\cdots=\check{\mathfrak{m}}_{k}.$
Definition 4.
The tuple $X=(X_{i})_{i\in I}$ are the refined leading terms of the canonical form of $\nabla$ with slopes $R$.
Lemma 5.
Fix refined leading terms $X=(X_{i})_{i\in I}$. For any $1\leq i\leq k$, let $i_{s}$ be the maximal element $i_{s}\in I$ with $i_{s}\leq i$. Then
(9) $\check{M}_{i}=\check{M}_{i_{s}}=C_{\check{G}}(X_{1},...,X_{i_{s}}).$
In particular, the Levi sequence $\check{M}_{i}$ is determined by the refined leading terms.
Proof.
First, $\check{M}_{1}$ is the centralizer of $X_{1}=D_{1}$. We have $D_{i}\in\check{\mathfrak{z}}_{1}$ and $\check{M}_{i}=\check{M}_{1}$ for $1\leq i<i_{2}$. Write $D_{i_{2}}=Z_{i_{2}}+X_{i_{2}}$ for the decomposition $\check{\mathfrak{z}}_{i_{2}}=\check{\mathfrak{z}}_{i_{2}-1}\oplus\check{\mathfrak{a}}_{i_{2}}$. Then an element of $\check{M}_{i_{2}-1}=\check{M}_{1}$ centralizes $D_{i_{2}}$ if and only if it centralizes $X_{i_{2}}$. Repeat the process determines all the Levi subgroups $\check{M}_{i}$ as in (9). ∎
Definition 6.
A tuple $X=(X_{i}\in\check{\mathfrak{a}}_{i})_{i\in I}$ is generic for the sequence $\check{M}_{i},1\leq i\leq k$ if (9) is satisfied.
Since the irregular part of any canonical form of a formal connection is unique up to $\check{G}(\mathbb{C})$-conjugation, the set $R$, the sequence of Levi subgroups $\check{M}_{i}$, and the refined leading terms $(X_{i})_{i}$ are all uniquely determined by the isomorphic class of the formal connection up to conjugation.
The following property follows immediately from definition.
Lemma 7.
If the refined leading terms of $\nabla$ are generic for $\check{M}_{i}$ with slopes $r_{i}$, $i\in I$, then the slopes of the adjoint connection $\nabla^{\mathrm{Ad}}$ consist of $\dim\check{\mathfrak{m}}_{k+1}=\mathrm{rk}\check{\mathfrak{g}}$ copies of 00 and $|\Phi(\check{M}_{i_{j}})-\Phi(\check{M}_{i_{i+1}})|$ copies of $r_{i_{j}}$, $i_{j}\in I$.
2.2.2.
Recall there exists $\theta\in\check{G}$ satisfying $\theta^{b}=1$, $br_{i}\in\mathbb{Z}$, $\theta(D_{i})=\exp(-2\pi\sqrt{-1}r_{i})D_{i}$. The torsion inner automorphism $\mathrm{Ad}_{\theta}$ defines a grading
(10) $\check{\mathfrak{g}}=\bigoplus_{i\in\mathbb{Z}/b\mathbb{Z}}\check{\mathfrak{g}}_{i}.$
Since $D_{i}$’s are $\theta$-eigenvectors, $\theta$ acts on $\check{M}_{i}$, $\check{\mathfrak{m}}_{i}$, $\check{\mathfrak{z}}_{i}$, $\check{\mathfrak{m}}_{i}^{\mathrm{der}}$, $\check{\mathfrak{a}}_{i}$. For each of these Lie algebras $\mathfrak{c}$, denote $\mathfrak{c}_{j}=\mathfrak{c}\cap\check{\mathfrak{g}}_{j}$. Let $m$ be the minimal common denominator of all the $r_{i}$, so that $m|b$. Define
(11) $\mathrm{Loc}(R,\theta)=\mathrm{d}+(\sum_{i\in I}\sum_{-mr_{i}\leq j\leq-1}\check{\mathfrak{a}}_{i,j}t^{-\frac{j}{m}}+\check{\mathfrak{m}}_{k})\frac{\mathrm{d}t}{t},$
and
(12) $\mathrm{Loc}(X,R,\theta)=\mathrm{d}+\sum_{i\in I}(X_{i}t^{-r_{i}}+\sum_{-mr_{i}+1\leq j\leq-1}\check{\mathfrak{a}}_{i,j}t^{-\frac{j}{m}}+\check{\mathfrak{m}}_{k})\frac{\mathrm{d}t}{t}.$
Lemma 8.
Two equations with the same refined leading terms $\nabla_{1},\nabla_{2}\in\mathrm{Loc}(X,I,R,\theta)$ give isomorphic formal connections only if these equations (12) have identical irregular parts.
Proof.
Write the canonical forms of $\nabla_{1},\nabla_{2}$ as in (8):
$\nabla_{j}=\mathrm{d}+(D_{1,j}t^{-r_{1}}+\cdots+D_{k,j}t^{-r_{k}}+D_{k+1,j})\frac{\mathrm{d}t}{t},\quad j=1,2,$
Suppose $\nabla_{1},\nabla_{2}$ are isomorphic. Then there exists $g\in\check{G}$ such that $Ad_{g}(D_{i,1})=D_{i,2}$ for $r_{i}\neq 0$. Since $D_{1}=X_{1}$, $g\in C_{\check{G}}(X_{1})=\check{M}_{1}$. Thus $\mathrm{Ad}_{g}D_{i,1}=D_{i,1}=D_{i,2}$ for $1\leq i<i_{2}$. Write $D_{i_{2},j}=Z_{i_{2},j}+X_{i_{2}}$, $j=1,2$. Note $g\in\check{M}_{1}=\check{M}_{i_{2}-1}$ centralizes $Z_{i_{2},j}\in\check{\mathfrak{z}}_{i_{2}-1}$. Thus $Ad_{g}(D_{i_{2},1})=D_{i_{2},2}$ implies $Z_{i_{2},1}+\mathrm{Ad}_{g}X_{i_{2}}=Z_{i_{2},2}+X_{i_{2}}$. Note that $\check{M}_{i_{2}-1}$ acts on $\check{\mathfrak{m}}_{i_{2}-1}^{\mathrm{der}}$, so that $\mathrm{Ad}_{g}X_{i_{2}}-X_{i_{2}}=Z_{i_{2},2}-Z_{i_{2},1}\in\check{\mathfrak{m}}_{i_{2}-1}^{\mathrm{der}}\cap\check{\mathfrak{z}}_{i_{2}-1}=0$. We obtain $D_{i_{2},1}=D_{i_{2},2}$ and $g\in\check{M}_{i_{2}}$. Repeat the procedure, we obtain $g\in\check{M}_{k}$ and $D_{i,1}=D_{i,2}$ for all $i\leq k$. ∎
2.2.3.
Alternatively, we can characterize Levi subgroups $\check{M}_{i}$ as follows. Let $I_{\mathrm{diff}}$ be the differential Galois group of $D^{\times}$, so that each formal connection $\nabla$ corresponds to its monodromy representation $\rho_{\nabla}$ of $I_{\mathrm{diff}}$, each formal $\check{G}$-connection corresponds to a group homomorphism $\rho_{\nabla}:I_{\mathrm{diff}}\rightarrow\check{G}$. Let $I_{\mathrm{diff}}^{r}$(resp. $I_{\mathrm{diff}}^{r+}$) be the intersection of the kernels of those representations of $I_{\mathrm{diff}}$ corresponding to formal connections with slopes $<r$(resp. $\leq r$). Here $I_{\mathrm{diff}}^{0}=I_{\mathrm{diff}}$. For $r>0$(resp. $r\geq 0$), $\rho_{\nabla}(I_{\mathrm{diff}}^{r})$(resp. $\rho_{\nabla}(I_{\mathrm{diff}}^{r+})$) is a connected torus, thus can be conjugated into the maximal torus $\check{H}$.
Lemma 9.
Let $\rho_{\nabla}:I_{\mathrm{diff}}\rightarrow\check{G}$ be the monodromy representation of a formal connection $\nabla$. For $r>0$, denote root subsystem
$\Phi_{r}=\{\alpha\in\Phi(\check{G},\check{H})\mid\alpha(\rho_{\nabla}(I_{\mathrm{diff}}^{r}))=1\}.$
Then $R$ is the set of $r$ where $\Phi_{r}$ jumps, and $\Phi_{r_{i}}$, $r_{i}\in R$ is the root system of $\check{M}_{i}$.
Proof.
This is clear from the observation that the restricted representation $\rho_{\nabla}:I_{\mathrm{diff}}^{r}\rightarrow\check{G}$ is determined only by the coefficients $D_{i}$, $r_{i}\geq r$ in the canonical form (8). ∎
We will see that the components $(X_{i_{s}}t^{-r_{i}}+\sum_{\ell=1}^{s-1}\sum_{j=-mr_{i_{s-1}}+1}^{-mr_{i_{s}}}Y_{i_{\ell},j}t^{\frac{j}{m}})\frac{\mathrm{d}t}{t}\in\check{\mathfrak{z}}_{i}(\!(t)\!)\frac{\mathrm{d}t}{t}$, $Y_{i_{\ell},j}\in\mathfrak{a}_{i_{\ell},j}$ $i_{s}\in I$ should be dual to the Howe factorization [KalethaRegular, Definition 3.6.2], see discussion in §3.14.1.
2.3. Oper canonical forms
A formal connection can always be represented in a non-unique way by equations called the oper canonical forms, which we review.
A $\check{G}$-oper over a space is a $\check{G}$-bundle with a connection and a Borel reduction satisfying a strong transversality condition, c.f. [BD, FGLocal]. Denote by $\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})$ the scheme of opers over $D^{\times}$. Let $E_{-\alpha}\in\check{\mathfrak{g}}_{-\alpha}$ be a basis of the negative simple root subspace, $\alpha\in\Delta_{\check{G}}$ a simple root. Then $p_{-1}=\sum_{\alpha\in\Delta_{\check{G}}}E_{-\alpha}$ is a regular nilpotent element. Let $\{p_{-1},2\check{\rho}_{\check{G}},p_{1}\}$ is the unique principal $\mathfrak{sl}_{2}$-triple, and $p_{i}$ be a homogeneous basis of $\check{\mathfrak{g}}^{p_{1}}$ with weight $d_{i}-1$ under the action of $\check{\rho}$, $1\leq i\leq n$. Any $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})$ can be uniquely represented by an oper canonical form as follows:
(13) $\nabla_{\chi}=\mathrm{d}+(p_{-1}+\sum_{i=1}^{n}v_{i}(t)p_{i})\mathrm{d}t,\quad v_{i}(t)=\sum_{j}v_{ij}t^{-j-1}\in\mathbb{C}(\!(t)\!),\ v_{ij}\in\mathbb{C}.$
In particular, we have isomorphism
$\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})\simeq\mathbb{C}(\!(t)\!)^{n},\qquad\nabla\mapsto(v_{1}(t),...,v_{n}(t)).$
For an oper $\chi$ with canonical form (13), its slope is defined to be
(14) $s(\chi)=\sup\{0,\sup_{1\leq i\leq n}\{\frac{-\mathrm{ord}_{t}v_{i}(t)}{d_{i}}-1\}\}$
where we let $\mathrm{ord}_{t}(0)=+\infty$.
We have a forgetful functor from opers to the underlying local system forgetting the Borel reduction:
(15) $p:\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})\rightarrow\mathrm{Loc}_{\check{G}}(D^{\times}).$
Equivalently, $p$ sends an oper canonical form $\nabla_{\chi}$ to its $G(\!(t)\!)$-gauge equivalence class. It is proved in [FZOper] that $p$ is surjective, but it has infinite dimensional fibers.
By [CK, Proposition 1], the slope of an oper coincides with the slope of its underlying connection, so that $s(\chi)$ is constant on the fibers of $p$.
2.4. Relationship between connection canonical forms and oper canonical forms
2.4.1.
We describe another way deducing the sequence $r_{i_{1}},...,r_{i_{k}}$ and Levi subgroups $\check{M}_{i_{1}},...,\check{M}_{i_{k}}$ from a formal connection starting from its oper canonical form. Denote $r_{(j)}=r_{i_{j}}$, $\check{M}_{(j)}=\check{M}_{i_{k}}$, and similarly for $\check{\mathfrak{m}}_{(j)},\check{\mathfrak{z}}_{(j)},\check{\mathfrak{m}}_{(j)}^{\mathrm{der}}$.
Let $\nabla$ be an irregular formal connection. Let $\chi=\mathrm{d}+p_{-1}+\sum_{i,j}v_{i,j}t^{-j-1}p_{i}\mathrm{d}t$ be an oper canonical form of $\nabla$. Assume $\nabla$ has slope $r_{(1)}=r_{1}$.
By the oper slope formula (14), $\chi$ is of the form
(16) $\mathrm{d}+(p_{-1}+\sum_{i=1}^{n}\sum_{j+1\leq d_{i}(r_{1}+1)}v_{ij}t^{-j-1}p_{i})\mathrm{d}t.$
Applying gauge transform by $\check{\rho}(t^{r_{1}+1})$ to the above equation, we obtain
(17) $\mathrm{d}+(t^{-r_{1}}p_{-1}+\sum_{i=1}^{n}\sum_{-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)}v_{ij}t^{-j+(r_{1}+1)(d_{i}-1)}p_{i}-(r_{1}+1)\check{\rho})\frac{\mathrm{d}t}{t}.$
Adopting the algorithm in [BV, Proposition 4.6], we can gauge transform the above into a canonical form by induction. We sketch the algorithm in this situation in the following.
First, by the slope formula, there exists $1\leq i\leq n$ and $j$ such that $j+1=d_{i}(r_{1}+1)$ and $v_{i,j}\neq 0$. Let
(18) $D_{r_{1}}=p_{-1}+\sum_{j+1=d_{i}(r_{1}+1)}v_{i,j}p_{i}.$
Let $D_{r_{1}}=D_{r_{1},s}+D_{r_{1},n}$ be Jordan decomposition and denote $X_{r_{1}}=D_{r_{1},s}$, $Y_{r_{1}}=D_{r_{1},n}$. Let $\check{\mathfrak{m}}_{1}=\mathfrak{z}_{\check{\mathfrak{g}}}(X_{r_{1}})$. Observe that $D_{r_{1}}$ is in the Kostant section, so that it is regular. Equivalently, $Y_{r_{1}}$ is regular nilpotent in $\check{\mathfrak{m}}_{1}$. Let $\mathcal{O}_{\overline{F}}$ be the integer ring of algebraic closure $\overline{F}$, $\mathcal{O}_{\overline{F}}^{+}$ its prime ideal so that $\mathcal{O}_{\overline{F}}/\mathcal{O}_{\overline{F}}^{+}\simeq\mathbb{C}$. Let $\check{G}(\mathcal{O}_{\overline{F}})_{0+}$ be the kernel of the reduction map $\check{G}(\mathcal{O}_{\overline{F}})\rightarrow\check{G}(\mathbb{C})$. Picking an $\mathfrak{sl}_{2}$-triple contaning $X_{r_{1}}$ and applying gauge transforms by elements in $\check{G}(\mathcal{O}_{\overline{F}})_{0+}$, we can transfer (17) into the form
(19) $\nabla_{1}\in\mathrm{d}+((X_{r_{1}}+Y_{r_{1}})t^{-r_{1}}+t^{-r_{1}}\check{\mathfrak{m}}_{r_{1}}(\mathcal{O}_{\overline{F}})_{0+})\frac{\mathrm{d}t}{t}\subset\mathrm{d}+\check{\mathfrak{m}}_{1}(\overline{F})\mathrm{d}t.$
We decompose $\nabla_{1}$ according to the Levi decomposition $\check{\mathfrak{m}}_{1}=\check{\mathfrak{z}}_{r_{1}}+\check{\mathfrak{m}}_{1}^{\mathrm{der}}$ and write $\nabla_{1}=\nabla_{\check{\mathfrak{z}}_{1}}+\nabla_{\check{\mathfrak{m}}_{1}^{\mathrm{der}}}$. Recall $\check{\mathfrak{z}}_{1}=\check{\mathfrak{a}}_{1}$. Let $\check{Z}_{1}\subset\check{H}$ be the subtorus with Lie algebra $\check{\mathfrak{z}}_{1}$. We can transfer $\nabla_{1}$ into canonical form using $\check{Z}_{1}(\mathcal{O}_{\overline{F}})_{0+}$, so we assume it already is. Here $X_{r_{1}}$ is the leading term of $\nabla_{\check{\mathfrak{z}}_{1}}$.
The $\check{M}_{1}^{\mathrm{der}}$-connection $\nabla_{\check{\mathfrak{m}}_{1}^{\mathrm{der}}}$ has regular nilpotent leading term $Y_{r_{1}}$. Take an $\mathfrak{sl}_{2}$-triple $\{e_{1},f_{1},h_{1}\}$ inside $\check{\mathfrak{m}}_{1}^{\mathrm{der}}$ where $f_{1}=Y_{r_{1}}$. Using gauge transform by $\check{M}_{1}^{\mathrm{der}}(\mathcal{O}_{\overline{F}})_{0+}$, we can change $\nabla_{\check{\mathfrak{m}}_{1}^{\mathrm{der}}}$ into the form
$\nabla^{\prime}=\mathrm{d}+(Y_{r_{1}}t^{-r_{1}}+t^{-r_{1}}(\check{\mathfrak{m}}_{1}^{\mathrm{der}})^{e_{1}}(\mathcal{O}_{\overline{F}})_{0+})\frac{\mathrm{d}t}{t}.$
Let $(\check{\mathfrak{m}}_{1}^{\mathrm{der}})^{e_{1}}=\bigoplus_{1\leq k\leq q}\mathbb{C}Z_{k}$ be a homogeneous basis, where $[h_{1},Z_{k}]=\lambda_{k}Z_{k}$ for integer $\lambda_{k}\geq 0$. This is exactly an oper canonical form for $\check{\mathfrak{m}}_{1}^{\mathrm{der}}$. Denote $\Lambda(Y_{r_{1}})=\sup_{k}(\frac{1}{2}\lambda_{k}+1)$. Write
$\nabla^{\prime}=\mathrm{d}+(Y_{r_{1}}t^{-r_{1}}+\sum_{r>-r_{1}}\sum_{1\leq k\leq q}a_{r,k}t^{r}Z_{k})\frac{\mathrm{d}t}{t},\qquad a_{r,k}\in\mathbb{C}.$
In the above $r$ sums over rational numbers such that $a_{r,k}\neq 0$. Define
(20) $\delta(\nabla^{\prime})=\inf\{\frac{r-r_{1}}{\frac{1}{2}\lambda_{k}+1}\mid 0<r+r_{1}<\Lambda(Y_{r_{1}})|r|,\ a_{r,k}\neq 0\}.$
If there is no such $a_{r,k}=0$, let $\delta(\nabla^{\prime})=\infty$.
If $\delta(\nabla^{\prime})\neq\infty$, we apply gauge transform by $t^{-\frac{\delta(\nabla^{\prime})}{2}h}$ to $\nabla^{\prime}$. Otherwise apply gauge transform by $t^{-\frac{r_{1}}{2}h}$ to $\nabla^{\prime}$. The resulting connection $\nabla^{\prime\prime}$ is either regular singular, or is of order $r_{(2)}=r_{1}-\delta(\nabla^{\prime})$ whose leading term is of the form
(21) $D_{r_{(2)}}=Y_{r_{1}}+\sum_{k,r}a_{r,k}Z_{k}.$
In $\nabla^{\prime\prime}$ becomes regular singular, we stop. Otherwise, since $Y_{r_{1}}$ is regular nilpotent in $\check{\mathfrak{m}}_{1}^{\mathrm{der}}$, $\{e_{1},f_{1}=Y_{r_{1}},h_{1}\}$ is a principal $\mathfrak{sl}_{2}$-triple of $\check{\mathfrak{m}}_{1}^{\mathrm{der}}$, so that (21) is contained in a Kostant section of $\check{\mathfrak{m}}_{1}^{\mathrm{der}}$. If $\nabla^{\prime\prime}$ is not regular singular, then $\delta(\nabla^{\prime})\neq\infty$, there exists $a_{r,k}\neq 0$ in (21), so that (21) is a non-nilpotent regular element. We can now just repeat the previous procedure: take Jordan decomposition $D_{r_{(2)}}=X_{r_{(2)}}+Y_{r_{(2)}}$, let $\check{\mathfrak{m}}_{(2)}=\check{\mathfrak{z}}_{\check{\mathfrak{m}}_{1}}(X_{r_{2}})$ with center $\check{\mathfrak{z}}_{(2)}$. Note that the center of $\check{\mathfrak{z}}_{\check{\mathfrak{m}}_{1}^{\mathrm{der}}}(X_{r_{2}})$ equals to $\check{\mathfrak{a}}_{(2)}=\check{\mathfrak{z}}_{(2)}\cap\check{\mathfrak{m}}_{1}^{\mathrm{der}}$. Denote the canonical form of the direct summand of $\nabla^{\prime\prime}$ inside $\check{\mathfrak{a}}_{(2)}$ by $\nabla_{\check{\mathfrak{a}}_{(2)}}$.
This procedure produces a sequence of slopes $r_{(1)}>r_{(2)}>\cdots>r_{(p)}$ and Levi subalgebras $\check{\mathfrak{m}}_{1}=\check{\mathfrak{m}}_{(1)}\supset\check{\mathfrak{m}}_{(2)}\supset\cdots\supset\check{\mathfrak{m}}_{(p)}$. We conclude that the irregular part of the canonical form of $\nabla$ is
(22) $\bigoplus_{s=1}^{p}\nabla_{\check{\mathfrak{a}}_{(s)}}.$
2.4.2.
Denote
(23) $\mathrm{Op}_{\check{\mathfrak{g}}}(X,R,\theta)=p^{-1}(\mathrm{Loc}(X,R,\theta)).$
Then the canonical form of any such oper would induce slopes $r_{(1)},...,r_{(p)}$ and Levi subgroups $\check{M}_{(1)},...,\check{M}_{(p)}$ as in §2.4.1.
Proposition 10.
  • (i)
    The oper canonical form (13) of an oper $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(X,R,\theta)$ satisfies $v_{i,j}=0$ for $-j+(d_{i}-1)(r_{1}+1)<-r_{1}$.
  • (ii)
    Denote the Coxeter number of $\check{M}_{(s)}$ by $h_{s}$, which is set to be $1$ if $\check{M}_{(s)}$ is a torus. The direct summand $\nabla_{\check{\mathfrak{a}}_{(s)}}$ of the canonical form of $\chi$ in (i) depends only on $v_{i,j}$ for $-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)\leq\sum_{a=1}^{s-1}(h_{a}-1)(r_{(a)}-r_{(a+1)})$. In particular, the irregular part of the canonical form of $\chi$ depends only on $v_{i,j}$ for $-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)\leq\sum_{a=1}^{p-1}(h_{a}-1)(r_{(a)}-r_{(a+1)})$. Moreover, the regular part of the canonical form, therefore the whole canonical form, depends only on $v_{i,j}$ for $-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)\leq\sum_{a=1}^{p-1}(h_{a}-1)(r_{(a)}-r_{(a+1)})+(h_{p}-1)r_{(p)}$.
Proof.
Part (i) follows from the oper slope formula (14).
For part (ii), in the algorithm §2.4.1, beginning from equation (17), coefficients $v_{i,j}$ with $-j+d_{i}(r_{1}+1)>0$ get involved in the canonical form only when we apply gauge transform by $t^{-\frac{\delta(\nabla^{\prime})}{2}h}$. Precisely, coefficients of $t^{q}\mathrm{d}t$ with $q\geq 0$ becomes coefficients of $t^{q-(d_{i,(a)}-1)(r_{(a)}-r_{(a+1)})}\mathrm{d}t$, where $d_{i,(a)}$ is a fundamental degree of $\check{M}_{(a)}$ which is at most $h_{a}$. The irregular part of part (ii) follows immediately. The dependence of the regular part of canonical form follows from [BV, Proposition 4.6.(a)]. ∎
Remark 11.
Proposition 10.(ii) refines the bound given by [BV, Theorem 9.7] applied to (17).
3. Refined leading terms of tame elliptic pairs
We will refine Conjecture 1 for a large family of regular supercuspidal representations that includes the toral supercuspidal representations. In this section we review their construction and define their refined leading terms.
3.1. Recollection on certain regular supercuspidal representations
We first review the construction of regular supercuspidal representations after [KalethaRegular].
3.1.1.
Let $S\subset G_{F}$ be an elliptic maximal torus over $F=\mathbb{C}(\!(t)\!)$ that splits over $E=\mathbb{C}(\!(u)\!)$, $u=t^{\frac{1}{m}}$. Explicitly, fix a maximal torus $H\subset G$, let $w_{m}\in W(G_{E},H_{E})$ be an elliptic element of order $m$ in the Weyl group, $\sigma:u\mapsto\zeta_{m}u$ a generator of $\mathrm{Gal}(E/F)$. We assume $S\simeq(\mathrm{Res}_{E/F}(H_{E}))^{w_{m}\times\sigma^{-1}}$, so that $S_{E}\simeq H_{E}$. Let $\phi:S\rightarrow\mathbb{C}^{\times}$ be a character of depth $\nu$ satisfying conditions in [KalethaRegular, Definition 3.7.5]. Such $(S,\phi)$ is called a tame elliptic pair. We further assume that
(24) $\{\alpha\in\Phi(G,S)\mid\phi(N_{E/F}(\check{\alpha}(E^{\times}_{0+})))=1\}=\varnothing.$
where $E^{\times}_{0+}=1+u\mathbb{C}[\![u]\!]$. This condition is to avoid the appearance of the depth zero supercuspidal representation in Yu’s datum. It corresponds to those L-parameters such that the centralizers of the images of the wild inertia group are tori.
We associate a Yu datum to $(S,\phi)$ as in [KalethaRegular, Proposition 3.7.8]. We slightly modify the notations to simplify them in our setting and to be more consistent with the notations on the spectral side. First, $S$ has a unique point $x$ in its Bruhat-Tits building, giving a point in the building of $G$. We can then define a sequence of root subsystems
$\Phi_{r}=\{\alpha\in\Phi(G,S)\mid\phi(N_{E/F}(\check{\alpha}(E^{\times}_{r})))=1\},$
where $E_{r}^{\times}=1+u^{\lceil mr\rceil}\mathbb{C}[\![u]\!]$. By Lemma 3.6.1 of loc. cit., $\Phi_{r}$ is a Levi subsystem. Let $R=\{r_{(1)},r_{(2)},...,r_{(p)}\}$ be real numbers $r_{(1)}>r_{(2)}>\cdots>r_{(p)}>0$ where $\Phi_{r}$ takes a jump. Note $r_{(1)}=\nu$. Let $G_{i}^{\prime}\subset G(E)$ be the standard Levi subgroup containing $S_{E}$ with root system $\Phi_{r_{(i)}}$. Then $G_{i}=G(F)\cap G_{i}^{\prime}$ define twisted Levi subgroups of $G(F)$:
$G=G_{0}\supset G_{1}\supset G_{2}\supset\cdots\supset G_{p}=S.$
By [KalethaRegular, Proposition 3.6.7], character $\phi$ has a Howe factorization as in Definition 3.6.2 of loc. cit.. It consists of a sequence of characters $\phi_{i}:G_{i}\rightarrow\mathbb{C}^{\times}$, $1\leq i\leq p$ satisfying
(25) $\phi=\prod_{i=1}^{p}\phi_{i}|_{S(F)}.$
Here two of $\phi_{i}$’s in loc. cit. are trivial because of our assumptions (24) and $\nu=r_{(1)}$. For $1\leq i\leq p$, $\phi_{i}$ is assumed to be trivial on $G_{i+1,sc}(F)$ where $G_{i+1,sc}$ is the simply-connected cover of the derived group of $G_{i+1}$, and $\phi_{i}$ has depth $r_{(i)}$. Moreover, $\phi_{i}$ needs to be $G_{i}$-generic in the sense of [Yu, §9], which we recall below.
Let $Z_{i}=Z(G_{i})$, with Lie algebra $\mathfrak{z}_{i}\subset\mathfrak{g}_{i}$. The dual $\mathfrak{z}_{i}^{*}$ can be naturally identified via restriction with the subspace of $\mathfrak{g}_{i}^{*}$ fixed under coadjoint action of $G_{i}$. Also, $\mathfrak{g}_{i}^{*}$ can be identified via restriction with the subspace of $\mathfrak{g}_{i-1}^{*}$ fixed by $Z_{i}$. Thus $\mathfrak{z}_{i}^{*}\subset\mathfrak{g}_{i}^{*}\subset\mathfrak{g}_{i-1}^{*}$.
Denote by $\mathfrak{z}_{i,-r}^{*}$, $\mathfrak{g}_{i,-r}^{*}$, $\mathfrak{g}_{i-1,-r}^{*}$ the Moy-Prasad filtration on the dual spaces. For example, $\mathfrak{g}_{i,-r}^{*}$ is the subspace of $X\in\mathfrak{g}_{i}^{*}$ satisfying $\mathrm{ord}_{t}(X(\mathfrak{g}_{i,r+}))>0$. We say $\phi_{i}:G_{i}\rightarrow\mathbb{C}^{\times}$ is $G_{i-1}$-generic of depth $r_{(i)}$ if $\phi_{i}$ is trivial on $G_{i,r_{(i)}+}$, non-trivial on $G_{i,r_{(i)}}$, and its differential lands inside $\mathrm{d}\phi_{i}\in\mathfrak{z}_{i,-r_{(i)}}^{*}$ such that
(26) $\mathrm{ord}(\mathrm{d}\phi_{i}(H_{\alpha}))=-r_{(i)},\qquad\forall\ \alpha\in\Phi(G_{i-1},S)\backslash\Phi(G_{i},S).$
Here for a root $\alpha$ with coroot $\check{\alpha}:\mathbb{G}_{m}\rightarrow S$, $H_{\alpha}=d\check{\alpha}(1)\in\mathfrak{g}_{i-1}$. This is condition GE1 in [Yu, §8], and by Lemma 8.1 of loc. cit. GE1 implies GE2 of [Yu, §8] in our case.
3.1.2.
Denote
(27) $(G_{i})_{x,r,r^{\prime}}=G(F)\cap\langle T(E)_{r},U_{\alpha}(E)_{x,r},U_{\beta}(E)_{x,r^{\prime}}\mid\alpha\in\Phi(G_{i},S),\beta\in\Phi(G_{i},S)\backslash\Phi(G_{i+1},S)\rangle.$
Let $(G_{i})_{x,r,r^{\prime}+}=\cup_{s>r^{\prime}}(G_{i})_{x,r,s}$. The quotient $V_{i}=(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2}/(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2+}$ is commutative. Extend $\phi_{i+1}$ to $G_{i}$ via projection with respect to root subspaces. The pairing $\langle a,b\rangle=\phi_{i+1}(aba^{-1}b^{-1})$ defines a symplectic form on $V_{i}$. Choose a Lagrangian subspace $\mathfrak{m}_{i}\subset V_{i}$. Denote by $(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2,\mathfrak{m}_{i}}$ the preimage of $\mathfrak{m}_{i}$ in $(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2}$  [^3] In the usual Yu’s construction, instead of making a Lagrangian, Heisenberg representation and Weil representation are used. We make this change so that the eventual representation is induced from a character. The same change is made in [BBMY, JKY]. Denote
(28) $J=S_{0+}\prod_{i=0}^{p-1}(G_{i})_{x,r_{(i+1)},r_{(i+1)}/2,\mathfrak{m}_{i}},$
where we set $r_{(p+1)}=0$. Note that by our assumption, $(G_{p})_{x}=S_{0}\subset J$. The quotient $S_{0}/S_{0+}$ is the fixed points of a maximal torus by a regular elliptic element of Weyl group, which is a nontrivial finite group in general. Here we use $S_{0+}$ instead of $(G_{p})_{x}$ so that $J$ is connected and pro-unipotent.
Consider character $\tilde{\phi}=\prod_{j=1}^{p}\phi_{i}$ on $J$. Here $\phi_{i}$ are extended to $J$ as follows.
On $(G_{j})_{x,r_{(j+1)},r_{(j+1)}/2,\mathfrak{m}_{j}}\subset(G_{j})_{x,r_{(j+1)},r_{(j+1)}/2}$ where $i\leq j$, $G_{i}\supset G_{j}$, $\phi_{i}$ is restricted from $G_{i}$.
For $i>j$ where $G_{i}\subset G_{j}$, the character $\phi_{i}$ is first extended to $G_{x,r_{(i+1)}/2+}$ via projection
$G_{x,r_{(i+1)}/2+}/G_{x,r_{(i+1)}+}\simeq$ $\mathfrak{g}_{x,r_{(i+1)}/2+}/\mathfrak{g}_{x,r_{(i+1)}+}$
$\rightarrow$ $(\mathfrak{g}_{i})_{x,r_{(i+1)}/2+}/(\mathfrak{g}_{i})_{x,r_{(i+1)}+}\simeq(G_{i})_{x,r_{(i+1)}/2+}/(G_{i})_{x,r_{(i+1)}+},$
where the projection in the middle is via root subspace decomposition
$\mathfrak{g}(E)=\mathfrak{g}_{i}(E)\oplus\bigoplus_{\alpha\in\Phi(G,S)-\Phi(G_{i},S)}\mathfrak{g}_{\alpha}(E).$
Then restrict the above extension of $\phi_{i}$ to $(G_{j})_{x,r_{(j+1)},r_{(j+1)}/2,\mathfrak{m}_{j}}\subset(G_{j})_{x,r_{(j+1)}/2}\subset(G_{j})_{x,r_{(i+1)}/2+}\subset G_{x,r_{(i+1)}/2+}$.
3.2. Refined leading terms
Denote $\mathfrak{s}=\mathrm{Lie}(S)$, $\mathfrak{s}_{0}=\mathrm{Lie}(S_{0})=\mathrm{Lie}(S_{0+})=\mathfrak{s}_{0+}$. The character $\phi$ induces a vector $\mathrm{d}\phi\in\mathfrak{s}^{*}$. As on the spectral side, denote
$\mathfrak{a}_{i}:=\mathfrak{z}_{i}\cap\mathfrak{g}_{i-1}^{\mathrm{der}},\quad\mathfrak{z}_{i}=\bigoplus_{1\leq j\leq i}\mathfrak{a}_{j},\quad\mathfrak{z}_{i,-r_{(i)}}^{*}=\bigoplus_{1\leq j\leq i}\mathfrak{a}_{j,-r_{(j)}}^{*}.$
Note $\mathfrak{a}_{1}=\mathfrak{z}_{1}$, $\mathfrak{z}_{p}=\mathfrak{s}$. Here $\mathrm{d}\phi|_{\mathfrak{a}_{i}}$ has depth $r_{(i)}$.
The point $x$ in the building of $S$ induces a grading on $\mathfrak{s}_{E}$ of order $m$, which induces a grading on $\mathfrak{s}^{*}$, allowing us to define a section of $\mathfrak{s}_{r}^{*}/\mathfrak{s}_{r+}^{*}$ into $\mathfrak{s}^{*}$. In particular, $\mathfrak{a}_{j,-r_{(j)}}^{*}$ has an induced grading with a summand isomorphic to $\mathfrak{a}_{j,-r_{(j)}}^{*}/\mathfrak{a}_{j,-r_{(j)}+}^{*}$. Denote by $\phi_{r_{(i)}}$ the projection of $\phi$ first to $\mathfrak{a}_{j,-r_{(j)}}^{*}$ then to this summand isomorphic to $\mathfrak{a}_{j,-r_{(j)}}^{*}/\mathfrak{a}_{j,-r_{(j)}+}^{*}$.
We call tuple $\{\phi_{(1)},...,\phi_{(p)}\}$ the refined leading terms of $\phi$.
4. A local geometric Langlands for some regular supercuspidal representations
We formulate a refinement of Conjecture 1 for those regular supercuspidal representations defined in §3.
4.1. Dual refined leading terms
Take a tame elliptic pair $(S,\phi)$ satisfying extra assumptions in §3, such that the unique point $x$ in its building is in the closure of fundamental alcove. As in §3, we associate to $(S,\phi)$ the tuple $r_{(1)}>\cdots>r_{(p)}>0$ and twisted Levi subgroups $G_{1}\supset\cdots\supset G_{p}=S$. The denominator $m$ is an elliptic number of the Weyl group $W(G,S)$. We assume $mr_{(j)}\in\mathbb{Z}$.
Let $R=\{r_{(1)},...,r_{(p)}\}$. Fix a maximal torus $\check{H}\subset\check{G}$ and isomorphism with dual torus $\check{H}\simeq\check{S}$, as well as their Lie algebras $\check{\mathfrak{h}}\simeq\check{\mathfrak{s}}$. Let $E=\mathbb{C}(\!(u)\!)$, $u=t^{\frac{1}{m}}$, and generator $\sigma(u)=\zeta_{m}u$ of Galois group $\mathrm{Gal}(E/F)$. There is elliptic element $w_{m}\in W$ of order $m$ such that the isomorphism $\check{\mathfrak{h}}\otimes_{\mathbb{C}}E\simeq\mathfrak{s}^{*}\otimes_{F}E$ is compatible with the action of $w_{m}\times\sigma^{-1}$ on the left and the Galois action of $\sigma$ on the right. Consider the grading defined by $\mathrm{Ad}_{w_{m}}$:
$\check{\mathfrak{h}}=\bigoplus_{i\in\mathbb{Z}/m\mathbb{Z}}\check{\mathfrak{h}}_{i}.$
Then we have isomorphism
$\check{\mathfrak{h}}_{-mr_{(j)}}\simeq\mathfrak{s}_{-r_{(j)}}^{*}/\mathfrak{s}_{-r_{(j)}+}^{*}$
for any $1\leq j\leq p$.
Let $\check{M}_{(j)}\subset\check{G}$ be the standard Levi subgroup containing $\check{H}$ that is dual to $G_{j}$, with Lie algebra $\check{\mathfrak{m}}_{(j)}$. By letting $\check{M}_{i_{j}}=\check{M}_{(j)}$, we can define Lie algebras $\check{\mathfrak{a}}_{(j)}$, $\check{\mathfrak{z}}_{(j)}\subset\check{\mathfrak{h}}$ from $\check{\mathfrak{m}}_{(j)}$ as in §2. We have isomorphism $\check{\mathfrak{z}}_{(j)}\otimes_{\mathbb{C}}E\simeq\mathfrak{z}_{j}(E)^{*}$. Here $\mathfrak{z}_{j}^{*}$ can be identified with the subspace of $\mathfrak{g}_{j}^{*}$ fixed by $G_{j}$, which embeds into $\mathfrak{s}_{j}^{*}$ via restriction. Moreover, for $i<j$, where $\mathfrak{g}_{i}\supset\mathfrak{g}_{j}$ and $\mathfrak{z}_{i}\subset\mathfrak{z}_{j}$, the image of $\mathfrak{z}_{i}^{*}\hookrightarrow\mathfrak{g}_{i}^{*}$ lands inside the image of $\mathfrak{z}_{j}^{*}\hookrightarrow\mathfrak{g}_{j}^{*}$ under the restriction map $\mathfrak{g}_{i}^{*}\twoheadrightarrow\mathfrak{g}_{j}^{*}$. This induces an embedding $\mathfrak{z}_{i}^{*}\hookrightarrow\mathfrak{z}_{j}^{*}$ because the composition with restriction $\mathfrak{z}_{i}^{*}\hookrightarrow\mathfrak{g}_{i}^{*}\twoheadrightarrow\mathfrak{g}_{j}^{*}\xrightarrow{\mathrm{res}_{\mathfrak{z}_{i}}}\mathfrak{z}_{i}^{*}$ is identity. Thus we have decomposition $\mathfrak{z}_{i}^{*}\simeq\mathfrak{z}_{i-1}^{*}\oplus\mathfrak{a}_{i}^{*}$, which induces natural isomorphism
(29) $\check{\mathfrak{a}}_{(j)}\otimes_{\mathbb{C}}E\simeq\mathfrak{a}_{j}(E)^{*},\qquad\check{\mathfrak{a}}_{(j)}\cap\check{\mathfrak{h}}_{-mr_{(j)}}\otimes_{\mathbb{C}}E\simeq\mathfrak{a}_{j,-r_{(j)}}^{*}/\mathfrak{a}_{j,-r_{(j)}+}^{*}.$
Let $\{X_{(j)}\in\mathfrak{a}_{(j)}\cap\check{\mathfrak{h}}_{-mr_{(j)}}\}$ be the images of refined leading terms $\{\phi_{r_{(j)}}\in\mathfrak{a}_{j,-r_{(j)}}^{*}/\mathfrak{a}_{j,-r_{(j)}+}^{*}\}$ of $\phi$ under (29).
4.2. The conjectural correspondence
4.2.1. A normal subgroup
Consider $\tau_{i}=\mathfrak{g}(F)\cap(\bigoplus_{0\neq\alpha\in\Phi(G_{i},S)}\mathfrak{g}(E)_{\alpha})$, $10\leq i\leq p-$ 1. Define
(30) $\mathfrak{j}^{\prime}=\mathfrak{g}_{x,r_{(1)}+}+\sum_{i=0}^{p-1}((\tau_{i})_{r_{(i+1)}/2}\cap(\mathfrak{g}_{i})_{x,r_{(i+1)},r_{(i+1)}/2,\mathfrak{m}_{i}})+\sum_{i=2}^{p}(\mathfrak{a}_{i})_{r_{(i)}+}.$
Let $\mathfrak{j}^{+}$ be the ideal of $\mathfrak{j}$ generated by subspace $\mathfrak{j}^{\prime}$. Denote the preimage of $\mathfrak{j}^{+}/\mathfrak{g}_{x,r_{(1)}}\subset\mathfrak{j}/\mathfrak{g}_{x,r_{(1)}}\simeq J/G_{x,r_{(1)}}$ in $J$ by $J^{+}$.
Lemma 12.
The linear form $\tilde{\phi}$ is trivial on $\mathfrak{j}^{+}$.
Proof.
The linear form $\tilde{\phi}$ evaluated on $\mathfrak{j}^{\prime}$ is by first projecting to $\mathfrak{s}_{0}$ by quotienting out root subspaces, then applying $\phi$. The first step annihilates $\tau_{i}$. From the definition of $r_{(i)}$, the depth of $\phi$ on $\mathfrak{a}_{i}$ is $r_{(i)}$. Since $\tilde{\phi}$ is the differential of a character, thus a Lie algebra homomorphism, it is then trivial on the ideal generated by $\mathfrak{j}^{\prime}$. ∎
Remark 13.
We expect that one can choose Lagrangians $\mathfrak{m}_{i}\subset V_{i}$ appropriately so that $\mathfrak{j}^{\prime}$ is itself an ideal of $\mathfrak{j}$. See (50) and Lemma 28 for the case of $p=1$.
4.2.2.
Denote
(31) $\bar{\mathfrak{j}}:=\mathfrak{j}/\mathfrak{j}^{+}.$
Recall $\mathfrak{j}^{+}\subset\mathfrak{j}\subset\mathfrak{g}[\![t]\!]$, so that $\mathfrak{j}^{+}+\mathbb{C}\mathbf{1}\subset\mathfrak{j}+\mathbb{C}\mathbf{1}$ are subalgebras of $\widehat{\mathfrak{g}}$. Consider
(32) $\mathrm{Vac}_{\mathfrak{j}^{+}}:=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{j}^{+}+\mathbb{C}\mathbf{1}}\mathbb{C}=\widehat{U}(\mathfrak{g})/\widehat{U}(\mathfrak{g})(\mathfrak{j}^{+}+\mathbb{C}\mathbf{1})=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{j}+\mathbb{C}\mathbf{1}}U(\bar{\mathfrak{j}})=\widehat{U}(\mathfrak{g})\otimes_{U(\mathfrak{j}+\mathbb{C}\mathbf{1})}U(\bar{\mathfrak{j}}).$
We have
$\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}}\simeq\mathrm{Hom}_{\mathfrak{j}}(U(\bar{\mathfrak{j}}),\widehat{U}(\mathfrak{g})\otimes_{U(\mathfrak{j}+\mathbb{C}\mathbf{1})}U(\bar{\mathfrak{j}}))\simeq\mathrm{Vac}_{\mathfrak{j}^{+}}^{J^{+}}.$
Note that $\bar{\mathfrak{j}}\simeq\mathfrak{s}_{0}/\mathfrak{s}_{0}\cap\mathfrak{j}^{+}$ is abelian. We have $U(\bar{\mathfrak{j}})\hookrightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}}$. Denote $\mathfrak{Z}_{\mathfrak{j}^{+}}=\mathrm{Im}(\mathfrak{Z}\rightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}})$. Then $\mathrm{Op}_{\mathfrak{j}^{+}}:=\mathrm{Spec}\mathfrak{Z}_{\mathfrak{j}^{+}}$ defines a subspace of opers on $D^{\times}$.
Define
(33) $A:=U(\bar{\mathfrak{j}})\cap\mathfrak{Z}_{\mathfrak{j}^{+}}\subset\mathrm{End}(\mathrm{Vac}_{\mathfrak{j}^{+}}),$
which fits into commutative diagram
[…]
Conjecture 14.
  • (i)
    Given a sequence of twisted Levi subgroups $G_{1},...,G_{p}$ and a collection of refined leading terms $\phi_{(i)}:\mathfrak{a}_{i}\rightarrow\mathbb{C}$, $1\leq i\leq p$ as in §3. Composed with projection $\bar{\mathfrak{j}}\rightarrow\mathfrak{a}_{i}$, they define $\ker\phi_{(i)}\subset U(\bar{\mathfrak{j}})$. Denote $A^{\prime}=A\cap_{i}\ker\phi_{(i)}\subset A$. Let $X_{(i)}\in\check{\mathfrak{a}}_{i}$ be the dual refined leading term. Let $\mathrm{Loc}(X_{(i)},R,\theta)$ be the space of connections with such refined leading terms. Then
    (35) $\mathrm{Spec}\mathfrak{Z}_{\mathfrak{j}^{+}}/A^{\prime}\simeq p^{-1}(\mathrm{Loc}(X_{(i)},R,\theta))$
    as subspaces of $\mathrm{Op}_{\mathfrak{j}^{+}}$. In particular, for any connection underlying an oper in $\mathrm{Spec}\mathfrak{Z}_{\mathfrak{j}^{+}}/A^{\prime}$, the associated sequence of Levi subgroups $\check{M}_{1},...,\check{M}_{p}$ are dual to $G_{1},...,G_{p}$.
  • (ii)
    For any character $\tilde{\phi}$ of $J$ lifted from $(S,\phi)$ as in §3, let $\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}=\mathfrak{Z}_{\mathfrak{j}^{+}}/\ker(\tilde{\phi}|_{A})$, i.e. the central support of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}:=\mathrm{Vac}_{\mathfrak{j}^{+}}/\ker\tilde{\phi}$. Then $\mathrm{Spec}(\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}})=p^{-1}(\nabla)$ for a unique formal connection $\nabla$ whose associated slopes are $R=\{r_{(1)},...,r_{(p)}\}$, associated Levi subgroups $\check{M}_{(1)},...,\check{M}_{{(p)}}$ are dual to $G_{1},...,G_{p}$, and the refined leading terms $X_{(1)},...,X_{(p)}$ are dual to refined leading terms $\phi_{(1)},...,\phi_{(p)}$ of $\phi$.
  • (iii)
    For all the $J$ and refined leading terms defined from pair $(S,\phi)$ as in §3, the union of the corresponding local systems $\mathrm{Loc}(X_{(i)},R,\theta)$ are all the irreducible formal connections such that the centralizer of the image of wild inertia group $\rho_{\nabla}(I_{\mathrm{diff}}^{0+})$ is a maximal torus.
  • (iv)
    If two such pairs $(S,\phi)$ and $(S^{\prime},\phi^{\prime})$ are $G(F)$-conjugated, then the induced modules $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$, $\mathrm{Vac}_{\mathfrak{j}^{\prime},\tilde{\phi}^{\prime}}$ have the same central supports.
We will confirm the above conjecture for toral supercuspidal representations and irreducible isoclinic connections, see Remark 37.
Corollary 15.
Let $\nabla$ be a formal connection, $\chi$ an oper form of $\nabla$. Then $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ contains a nonzero $(J,\tilde{\phi})$-object for some $(J,\tilde{\phi})$ as in §3 if and only if $\nabla$ is an irreducible formal connection such that the centralizer of the image of wild inertia group is a maximal torus.
Proof.
If part: by (ii) and (iii) of Conjecture 14, there exists $(J,\tilde{\phi})$ such that any oper structure of $\nabla$ is in the central support of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$. We regard $\chi$ as a character of $\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}$. It suffices to show that the $(J,\tilde{\phi})$-equivariant object $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}/\ker\chi$ is nonzero. We show that its associated graded module is nonzero.
Denote by $(\ker\tilde{\phi})^{\perp}\subset\mathfrak{g}^{*}\otimes\omega_{F}$ the dual lattice of $\ker\tilde{\phi}$ and consider the local Hitchin map
$h^{cl}:(\ker\tilde{\phi})^{\perp}\rightarrow\mathrm{Hit}(D^{\times})$
as in [Zhu, §4.2, §4.3], see §6.4 for a review. Denote the Zariski closure of $h^{cl}((\ker\tilde{\phi})^{\perp})$ by $\mathrm{Hit}_{\tilde{\phi}}$. Then $\mathrm{gr}\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}=\mathrm{Fun}~{}\mathrm{Hit}_{\tilde{\phi}}$ and $\mathrm{gr}\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}=\mathrm{gr}(\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\ker\tilde{\phi}+\mathbb{C}\mathbf{1}}\mathbb{C})=\mathrm{Fun}~{}((\ker\tilde{\phi})^{\perp})$. Note that $0\in(\ker\tilde{\phi})^{\perp}$, so that $\mathrm{Hit}_{\tilde{\phi}}$ contains 00, the unique $\mathbb{G}_{m}$-fixed point of $\mathrm{Hit}(D^{\times})$. Let $H_{0}$ be the fiber over 00 of restriction of local Hitchin map $h^{cl}:(\ker\tilde{\phi})^{\perp}\rightarrow\mathrm{Hit}_{\tilde{\phi}}$. Then $H_{0}$, containing 00, is a nonempty closed subscheme of $(\ker\tilde{\phi})^{\perp}$.
Observe that $\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}$ can be identified with the image of $\mathfrak{Z}$ into $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$ via action on $1\in\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$, in particular is a subspace of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$. Also, $\mathrm{gr}(\ker\chi)$ is the maximal ideal of $\mathrm{Fun}~{}\mathrm{Hit}_{\tilde{\phi}}$ at 00. We obtain
$\mathrm{gr}(\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}/\ker\chi)\simeq\mathrm{gr}(\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}})/\mathrm{gr}(\ker\chi)\simeq\mathrm{Fun}~{}H_{0},$
which is nonzero. This implies that $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}/\ker\chi$ is nonzero.
Only if part: note that $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}\simeq\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\ker\tilde{\phi}+\mathbb{C}\mathbf{1}}\mathbb{C}$. By the same argument as the proof of [FrenkelBook, Lemma 10.3.2], there exists a nonzero object of $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ with a nonzero morphism from $\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\ker\tilde{\phi}+\mathbb{C}\mathbf{1}}\mathbb{C}$. Thus $\chi$ is in the central support of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$, so that its underlying connection is as in Conjecture 14.(iii). ∎
Remark 16.
  • (i)
    Part (iv) of Conjecture 14 is inspired by [KalethaRegular, Corollary 3.7.10].
  • (ii)
    We expect that the algebra $A$ is contained in the algebra generated by images of Segal-Sugawara operators $S_{i,j}\in\mathfrak{Z}$(see §6.2) satisfying the index condition $-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)\leq\sum_{a=1}^{p-1}(h_{a}-1)(r_{(a)}-r_{(a+1)})+(h_{p}-1)r_{(p)}$ as in Proposition 10. In fact, the proof of Proposition 29 for the toral supercuspidal case shows that the images of $S_{i,j}$ for $-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)<0$ are contained in $A$. However, unlike the toral supercuspidal case we will see in §7, in general $A$ is not expected to be freely generated by $S_{i,j}$: Lemma 22 indicates that if images of $S_{i,j}$ for $-r_{1}\leq-j+(d_{i}-1)(r_{1}+1)<0$ in $A$ are algebraically free, then $\check{M}_{(1)}$ should be a maximal torus.
  • (iii)
    If $A$ is generated by the images of those $S_{i,j}$ in (ii), then the underlying connection of any oper in the central support of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$ can be given by an oper canonical form whose coefficient $v_{i,j}$ for those $(i,j)$ in (ii) is given by $\tilde{\phi}(S_{i,j})$, and all other $v_{i,j}$ are zero. This provides an explicit recipe for the Langlands parameters. We will see in §7 that this is the case for toral supercuspidal representations
5. Isoclinic formal connections and opers
We study the relationship between oper canonical forms and connection canonical forms for isoclinic connections.
5.1. Isoclinic formal connections
5.1.1.
Following [JYDeligneSimpson], we call a formal connection $\nabla$ isoclinic if in its canonical form (8), the leading term $X_{r_{1}}$ is regular semisimple. For such connections, $R=\{r_{1}\}$, $p=1$, the sequence of Levi subgroups $\check{M}_{i_{s}}$ consists of only a single maximal torus centralizing $X_{r_{1}}$, and there are no higher leading terms beyond $X_{r_{1}}$. Clearly this notion is independent of the choice of canonical form and the gauge transformation. In particular, such $\nabla$ is irregular. The isoclinic formal connections are indeed the de Rham analog of the toral supercuspidal parameters [KalethaRegular, Definition 6.1.1].
For an irregular formal connection $\nabla$ with slope $\nu$, write $\nu=\frac{N}{m}$ where $(N,m)=1$. We collect some results on isoclinic formal connections from [JYDeligneSimpson]. Recall that $\check{\rho}:\mathbb{G}_{m}\rightarrow\check{T}$. Let $\zeta_{m}$ be a primitive $m$-th root of unity that is fixed once and for all. Let $\theta=\mathrm{Ad}_{\check{\rho}(\zeta_{m})}$ be an inner automorphism of $\check{\mathfrak{g}}$ that defines a grading
(36) $\check{\mathfrak{g}}=\bigoplus_{i\in\mathbb{Z}/m\mathbb{Z}}\check{\mathfrak{g}}_{i}.$
Lemma 17.
A formal connection $\nabla$ is isoclinic with slope $\nu=\frac{N}{m}>0$, $(N,m)=1$, if and only if $m$ is a regular number of $W$, and it is $G(\overline{F})$-gauge equivalent to a form as follows:
(37) $\mathrm{d}+(X_{-N}t^{-N/m}+\cdots+X_{-1}t^{-1/m}+X_{0})\frac{\mathrm{d}t}{t},$
where $X:=X_{-N}$ is regular semisimple, $X_{i}\in\mathfrak{z}_{\check{\mathfrak{g}}}(X)=:\check{\mathfrak{t}}_{X}$, $X_{i}\in\check{\mathfrak{g}}_{i}$. Moreover, if $m$ is elliptic, then $\nabla$ is irreducible.
Proof.
The if part follows from [BV, 9.8, Proposition].
For the only if part, first we know $m$ is a regular number from [JYDeligneSimpson, Lemma 2.2]. By [OY, Theorem 3.2.5], $x=\check{\rho}/m$ is $m$-regular in the sense of [JYDeligneSimpson, Definition 2.4]. By [JYDeligneSimpson, Lemma 2.10], $\nabla$ can be $G(\overline{F})$-gauge transformed to a form as (37) except $X_{0}$. Since $X$ is regular semisimple, we can use gauge transformation to make $X_{0}\in\check{\mathfrak{t}}_{X}$. Let $\check{T}_{X}=C_{\check{G}}(X)$, $W\simeq N_{\check{G}}(\check{T}_{X})/\check{T}_{X}$. Since $X$ is $\theta$-eigen, $\theta$ acts on $\check{\mathfrak{t}}_{X}$, where it acts as a regular element of order $m$ of $W$. By the proof of [JYDeligneSimpson, Lemma 2.2], there exists regular $w$ with the same action as $\theta$ on $X_{-N},...,X_{-1}$, and $w$ fixes $X_{0}$. Since $X\in\check{\mathfrak{t}}_{X}$ is a regular vector fixed by $w^{-1}\theta$, we have $\theta=w$ in $W$. Thus $X_{0}$ is fixed by $\theta$, i.e. $X_{0}\in\check{\mathfrak{g}}_{0}$.
Note that the differential Galois group of $\nabla$ contains the smallest torus whose Lie algebra contains $X$, together with a lift $w$ of an order $m$ regular element in the Weyl group $W(\check{G},\check{T}_{X})$. By the same proof as [CYTheta, Proposition 12.(v)], if $w$ is elliptic, $\nabla$ is irreducible. ∎
Lemma 18.
Two irreducible isoclinic formal connections are isomorphic, i.e. differ by a $\check{G}(F)$-gauge transform, if and only if they have isomorphic canonical forms, i.e. differ by a $\check{G}(\overline{F})$-gauge transform.
Proof.
Only the if part needs to be proved. Since we assumed $\check{G}$ simply, the proof is contained in the proof of [JYDeligneSimpson, Lemma 2.10] which we recall.
Let $\nabla,\nabla^{\prime}$ be irreducible isoclinic formal connections with isomorphic canonical forms. Let $\rho,\rho^{\prime}:I_{\mathrm{diff}}\rightarrow\check{G}$ be their monodromy representations. Then there exists a finite covering $D^{\times}_{a}=\mathrm{Spec}\mathbb{C}(\!(t^{1/a})\!)$ with differential Galois group $I_{\mathrm{diff}}^{a}\subset I_{\mathrm{diff}}$ such that the restrictions $\rho|_{I_{\mathrm{diff}}^{a}},\rho^{\prime}|_{I_{\mathrm{diff}}^{a}}$ can be $\check{G}$-conjugated to the same monodromy representation $\pi:I_{\mathrm{diff}}^{a}\rightarrow\check{T}$ landing inside a maximal torus $\check{T}$.
We may assume $\rho|_{I_{\mathrm{diff}}^{a}}=\rho^{\prime}|_{I_{\mathrm{diff}}^{a}}=\pi$. Note $I_{\mathrm{diff}}/I_{\mathrm{diff}}^{a}\simeq\mu_{a}$ is cyclic with a generator $\gamma$. By [JYDeligneSimpson, Lemma 2.2.(1)], $\rho(\gamma),\rho^{\prime}(\gamma)$ normalize $\check{T}$ and map to the same element $wT$ in the Weyl group, which must be elliptic by Lemma 17. For elliptic elements, their lifts to $\check{G}$ are all $\check{T}$-conjugated. Thus $\rho,\rho^{\prime}$ are $\check{T}$-conjugated. This proves the statement. ∎
5.1.2.
Denote by $\mathrm{Loc}(X,\nu)$ the isomorphism classes of irreducible isoclinic formal connections with slope $\nu$ and leading term $X$ in its canonical form. Any $\nabla\in\mathrm{Loc}(X,\nu)$ has a canonical form as in Lemma 17. Observe that its differential Galois group is contained in $\check{T}_{X}$ and a lift of order $m$ regular Weyl group element $w_{m}$. Thus for it to be irreducible, $w_{m}$ must have no nonzero fixed point on $\check{\mathfrak{t}}_{X}$, i.e. $w_{m}$ is elliptic. Equivalently, $m$ is a regular elliptic number of $W$, and the grading define by $\check{\rho}/m$ are exactly the inner stable gradings, see [RLYG] where such $m$ and stable gradings are classified.
Lemma 19.
The isomorphism classes of irreducible isoclinic connections $\mathrm{Loc}(X,\nu)$ with fixed leading term $X$ and slope $\nu$ are one-to-one corresponding to following equations:
(38) $\{\mathrm{d}+(X_{-N}t^{-N/m}+\cdots+X_{-1}t^{-1/m})\frac{\mathrm{d}t}{t}\mid X_{-N}=X,X_{i}\in\check{\mathfrak{t}}_{X}\cap\check{\mathfrak{g}}_{i}=:\check{\mathfrak{t}}_{X,i}\}.$
Proof.
By Lemma 18, isomorphism classes of irreducible isoclinic connections are in bijection with isomorphism classes of their canonical forms. If two canonical forms as in (37) with $X_{-N}=X$ are gauge equivalent, their irregular parts must differ by a gauge transform by some element $g\in\check{T}_{X}$, which acts as identity on such equations. Also, note that $X_{0}\in\check{\mathfrak{t}}_{X,0}=0$ as $w_{m}$ is elliptic. Thus such canonical form is unique. ∎
5.1.3.
We collect an observation for future use.
Lemma 20.
Assume $N\geq 1$. An irregular formal connection of the form
$\nabla=\mathrm{d}+\sum_{i=-N}^{\infty}X_{i}t^{i}\frac{\mathrm{d}t}{t},\quad X_{i}\in\check{\mathfrak{g}},$
with $X_{-N}$ regular semisimple has the same canonical form as $\nabla^{\prime}=\mathrm{d}+\sum_{i=-N}^{0}X_{i}t^{i}\frac{\mathrm{d}t}{t}$.
Proof.
Let $X=X_{-N}$, $\check{\mathfrak{t}}_{X}=\mathfrak{z}_{\check{\mathfrak{g}}}(X)$ be a Cartan subalgebra. Write $\check{\mathfrak{g}}=\check{\mathfrak{t}}_{X}\bigoplus_{\alpha\in\Phi(\check{\mathfrak{g}},\check{\mathfrak{t}}_{X})}\check{\mathfrak{g}}_{\alpha}$. Let $\pi_{\alpha}$ be the projection from $\check{\mathfrak{g}}$ to $\check{\mathfrak{g}}_{\alpha}$.
The first step in transforming $\nabla$ and $\nabla^{\prime}$ into canonical form is applying the gauge transforms to make them into $\mathrm{d}+\check{\mathfrak{t}}_{X}(\!(t)\!)\mathrm{d}t$. Explicitly, we gauge transofrm by convergent products of elements of the form $\exp(-\alpha(X)^{-1}\pi_{\alpha}(X_{i}^{\prime})t^{N+i})$ for $\alpha\neq 0$ and $i=-N+1,-N+2,...$, where $X_{-N+1}^{\prime}=X_{-N+1}$, and $X_{i}^{\prime}$ for $i>-N+1$ are determined by recursive relations. We obtain two formal connections of the form $\mathrm{d}+\sum_{i=-N}^{\infty}Y_{i}t^{i}\frac{\mathrm{d}t}{t}$ where $Y_{-N}=X$, all $Y_{i}\in\check{\mathfrak{t}}_{X}$, and $Y_{-N},...,Y_{0}$ are determined by $X_{-N},...,X_{0}$. The two connections only differ on their regular parts.
The second step is getting rid of the regular part. This is done by applying the gauge transform by $\exp(-\sum_{i\geq 1}t^{i}i^{-1}Y_{i})\in\check{T}_{X}(\!(t)\!)$, which does not change $Y_{i}$ for $i<0$. The two connections arrive at the same canonical form. ∎
Remark 21.
The above lemma is false for regular formal connections.
5.2. Oper canonical forms of irreducible isoclinic formal connections
Let $m$ be a regular elliptic number of $W$, $(N,m)=1$, $\nu=\frac{N}{m}$. We know from [JYDeligneSimpson, Lemma 2.5] that $\check{\mathfrak{g}}_{-N}$ contains regular semisimple elements. Let $X\in\check{\mathfrak{g}}_{-N}^{\mathrm{rs}}$, $\check{\mathfrak{t}}_{X}=\mathfrak{z}_{\check{\mathfrak{g}}}(X)$, $\check{T}_{X}=C_{\check{G}}(X)$. Since $X$ is $\theta$-eigen, $\theta$ acts on $\check{\mathfrak{t}}_{X}$, $\check{\mathfrak{t}}_{X}=\bigoplus_{i\in\mathbb{Z}/m\mathbb{Z}}\check{\mathfrak{t}}_{X,i}$.
Denote the space of opers whose underlying connection belong to $\mathrm{Loc}(X,\nu)$ by
(39) $\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu):=p^{-1}(\mathrm{Loc}(X,\nu)).$
Define
(40) $\ell_{i,j}:=(d_{i}-1)N+m(d_{i}-1-j)$
and
(41) $A_{\nu}=\{(i,j)\mid 1\leq i\leq n,\ -N+1\leq\ell_{i,j}\leq-1\},\quad\tilde{A}_{\nu}=\{(i,j)\mid 1\leq i\leq n,\ -N\leq\ell_{i,j}\leq-1\}.$
Lemma 22.
  • (i)
    The set $A_{\nu,\ell}:=\{(i,j)\in A_{\nu}\mid\ell=\ell_{i,j}\}$ has cardinality $\dim\check{\mathfrak{t}}_{X,\ell}$ for any $\ell$.
  • (ii)
    $\mathrm{Loc}(X,\nu)\simeq\mathbb{A}^{|A_{\nu}|}$.
Proof.
Denote $k_{\ell}:=|\{1\leq i\leq n\mid d_{i}-1\equiv\ell\mod m\}|$. Denote by $N^{-1}\in\mathbb{Z}/m\mathbb{Z}$ the inverse of $N\mod m$. We have
|$|\{(i,j)\in A_{\nu}\mid\ell=\ell_{i,j}\}|$
$=$ |$|\{1\leq i\leq n\mid\ell\equiv(d_{i}-1)N\mod m\}|$
$=$ |$|\{1\leq i\leq n\mid d_{i}-1\equiv N^{-1}\ell\mod m\}|$
$=$ $k_{N^{-1}\ell}.$
Let $\theta^{\prime}=\theta^{-N^{-1}}$, $\check{\mathfrak{g}}_{i}^{\prime}=\check{\mathfrak{g}}_{-N^{-1}i}$ the $\zeta_{m}^{i}$-eigenspace of $\theta^{\prime}$. Then $X\in\check{\mathfrak{g}}_{-N}=\check{\mathfrak{g}}_{1}^{\prime}$ is regular semisimple. Applying [Panyushev, Theorem 4.2.(i)] and its proof to $\theta^{\prime}$, we obtain
$k_{N^{-1}\ell}=\dim(\check{\mathfrak{t}}_{X}\cap\check{\mathfrak{g}}_{-N^{-1}\ell}^{\prime})=\dim(\check{\mathfrak{t}}_{X}\cap\check{\mathfrak{g}}_{\ell})=\dim(\check{\mathfrak{t}}_{X,\ell}).$
This proves part (i). Part (ii) follows from (i) immediately. ∎
By Lemma 19, we have isomorphism
$\mathrm{Loc}(X,\nu)\simeq\bigoplus_{-N<\ell\leq-1}\check{\mathfrak{t}}_{X,\ell}.$
Fix a basis $\{E_{a}^{\ell}\mid 1\leq a\leq\dim\check{\mathfrak{t}}_{X,\ell}\}$ of $\check{\mathfrak{t}}_{X,\ell}$. Let $e_{a}^{\ell}$ be the coordinate on $\mathrm{Loc}(X,\nu)$ for $E_{a}^{\ell}$.
Proposition 23.
  • (i)
    The space $\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)$ is isomorphic to the following affine space of oper canonical forms:
    $\{\mathrm{d}+(p_{-1}+\sum_{i=1}^{n}\sum_{\ell_{i,j}\geq-N}v_{i,j}t^{-j-1}p_{i})\mathrm{d}t\mid\ v_{i,j}\in\mathbb{C},\ p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}\in\check{G}\cdot X\}.$
  • (ii)
    In terms of the basis $v_{i,j}$ and $E_{a}^{\ell}$, the surjection $p:\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)\rightarrow\mathrm{Loc}(X,\nu)$ between two affine spaces is of the form
    (42) $e_{a}^{\ell}=f_{a,\ell}(v_{ij}\mid\ell_{i,j}<\ell)+\sum_{\ell_{i,j}=\ell}c_{(i,j),(a,\ell)}v_{i,j},$
    where $-N<\ell\leq-1$, $f_{a,\ell}$ is a polynomial in those $v_{ij}$ satisfying $\ell_{i,j}<\ell$, $c_{(i,j),(a,\ell)}\in\mathbb{C}$.
  • (iii)
    The map $p:\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)\rightarrow\mathrm{Loc}(X,\nu)$ factors as the projection forgetting $v_{i,j}$, $\ell_{i,j}\geq 0$ composed with an isomorphism
    (43) $\begin{split}&\{\mathrm{d}+(p_{-1}+\sum_{i=1}^{n}\sum_{-N\leq\ell_{i,j}\leq-1}v_{ij}t^{-j-1}p_{i})\mathrm{d}t\mid\ v_{i,j}\in\mathbb{C},\ p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}\in\check{G}\cdot X\}\\ &\simeq\mathrm{Loc}(X,\nu).\end{split}$
    In particular, the linear transformation $\phi_{\ell}:(v_{i,j})_{\ell_{i,j}=\ell}\mapsto(\sum_{\ell_{i,j}=\ell}c_{(i,j),(a,\ell)}v_{i,j})_{a}$ is invertible for any $-N<\ell\leq-1$ with $\dim\check{\mathfrak{t}}_{X,\ell}>0$.
Proof.
(i): Let $\nabla\in\mathrm{Loc}(X,\nu)$. Let $\nabla_{\chi}\in p^{-1}(\nabla)$ with oper canonical form (13). From the slope formula (14), $v_{i,j}=0$ for $\ell_{i,j}<-N$, and at least one of $v_{i,j}$ with $\ell_{i,j}=-N$ is nonzero.
Fix $u=t^{1/m}$. Applying $\mathrm{Ad}_{\check{\rho}(u^{N+m})}$ to $\nabla_{\chi}$, we obtain
(44) $\mathrm{d}+m[u^{-N}(p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i})+\sum_{i=1}^{n}\sum_{\ell_{i,j}>-N}v_{i,j}u^{\ell_{i,j}}p_{i}]\frac{\mathrm{d}u}{u}-(N+m)\check{\rho}\frac{\mathrm{d}u}{u}.$
Note that the leading term of $\nabla$ with respect to $u$ is $mX$, due to $\frac{\mathrm{d}t}{t}=m\frac{\mathrm{d}u}{u}$. By definition, the semisimple part of $p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}$ is conjugated to $X$. Since $X$ is regular semisimple, $p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}$ must also be. Thus $p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}\in\check{G}\cdot X$ and is the unique such element in the Kostant section. This proves (i).
(ii): We resume the computation of canonical form in part (i). By Lemma 20, we can drop $v_{i,j}$ for $\ell_{i,j}>0$. Also, since $m$ is elliptic, $\check{\mathfrak{t}}_{X,0}=0$, so that the canonical form has vanishing polar term. This proves the first assertion in part (iii).
To reduce the equation into a canonical form as in (38), we first conjugate by a fixed constant element $g\in\check{G}$ such that $\mathrm{Ad}_{g}(p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i})=X$. Next we apply the same algorithm as in the proof of Lemma 20 to project $v_{i,j}u^{\ell_{i,j}}p_{i}$ into $\check{\mathfrak{t}}_{X,\ell_{i,j}}u^{\ell_{i,j}}$ and eliminate the regular part. Part (ii) is clear from the calculation procedure.
(iii): We have already seen in the above discussion that $p$ factors through the projection to
$\{\mathrm{d}+(p_{-1}+\sum_{i=1}^{n}\sum_{-N\leq\ell\leq-1}v_{ij}t^{-j-1}p_{i})\mathrm{d}t\mid\ v_{i,j}\in\mathbb{C},\ p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}\in\check{G}\cdot X\},$
on which it is isomorphic to a surjective linear map from $\{v_{i,j}\mid-N<\ell_{i,j}\leq-1\}$ to $\{e_{a}^{\ell}\mid-N<\ell\leq 0\}$. By Lemma 22, this linear map must be invertible. This completes the proof of (iii). ∎
We can regard the above (43) as the ‘minimal’ oper canonical form of the formal connection.
Remark 24.
  • (i)
    Let $S$ be the centralizer of $\tilde{X}=p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}t^{-d_{i}-\frac{d_{i}N}{m}}p_{i}$. Then the connections in (43) are called $S$-toral of slope $\nu$ in [KSRigid].
  • (ii)
    When $m$ is not necessarily elliptic, we still have a slight variant of Proposition 23 with $\mathrm{Loc}(X,\nu)$ replaced by irregular types of such formal connections.
5.3. Global extension
We write down an extension to $\mathbb{P}^{1}-\{0\}$ of the isoclinic formal connections $\mathrm{Loc}(X,\nu)$.
Proposition 25.
Assume $\nu=\frac{N}{m}$. Let $\nabla$ be a $\check{G}$-connection over $\mathbb{G}_{m}$ with following equation:
(45) $\mathrm{d}+(t^{-2}p_{-1}+\sum_{i=1}^{n}\sum_{-N\leq\ell_{i,j}\leq-1}v_{ij}t^{-j+2d_{i}-3}p_{i})\mathrm{d}t,\quad p_{-1}+\sum_{m|d_{i}N}v_{i,d_{i}+\frac{d_{i}N}{m}-1}p_{i}\in\check{G}\cdot X.$
Denote by $\mathrm{Loc}(X,\nu,\mathbb{P}^{1}-\{0\})$ the isomorphism classes of such connections.
  • (i)
    $\nabla|_{D_{0}^{\times}}$ is isomorphic to the isoclinic formal connection
    $\mathrm{d}+(p_{-1}+\sum_{i=1}^{n}\sum_{-N\leq\ell_{i,j}\leq-1}v_{ij}t^{-j-1}p_{i})\mathrm{d}t.$
    The restriction to $D_{0}^{\times}$ induces an isomorphism $\mathrm{Loc}(X,\nu,\mathbb{P}^{1}-\{0\})\simeq\mathrm{Loc}(X,\nu)$ factoring through (43). In particular, every connection in $\mathrm{Loc}(X,\nu,\mathbb{P}^{1}-\{0\})$ is represented by a unique equation of the form (45).
  • (ii)
    When $\nu\geq 1$, $\nabla$ has trivial monodromy at $\infty$.
Proof.
(i): Replace $t=u^{m}$ and apply gauge transform by $\check{\rho}(u^{N-m})$ to (45). We obtain
$\mathrm{d}+m(p_{-1}u^{-N}+\sum_{i=1}^{n}\sum_{-N\leq\ell_{i,j}\leq-1}v_{ij}u^{\ell_{i,j}}p_{i})\frac{\mathrm{d}u}{u}-(N-m)\check{\rho}\frac{\mathrm{d}u}{u}.$
By the proof of Proposition 23.(ii), the above equation has the same canonical form as (44), thus is isomorphic to a connection as (43). Part (i) follows from Lemma 19 and Proposition 23.
(ii): Denote $s=t^{-1}$. We have
$\nabla|_{D_{\infty}^{\times}}\simeq\mathrm{d}-(p_{-1}+\sum_{i=1}^{n}\sum_{-N\leq\ell_{i,j}\leq-1}v_{ij}s^{j-2d_{i}+1})\mathrm{d}s.$
Note that
$\ell_{i,j}\leq-1\quad\Rightarrow\quad\frac{(d_{i}-1)(N-m)+1}{m}-1\leq j-2d_{i}+1.$
Since $N\geq m$, $j-2d_{i}+1>-1$, so that $\nabla|_{D_{\infty}^{\times}}$ is regular. ∎
Remark 26.
  • (i)
    In the above and the proof of surjectivity in (43), we did not essentially use the ellipticity of $m$. Thus we gain an explicit construction of a connection on $\mathbb{P}^{1}-\{0\}$ with arbitrary isoclinic irregular type at 00. In general, it is proved in [JYDeligneSimpson, Corollary 1.14.(ii)] that there exists a connection on $\mathbb{G}_{m}$ with arbitrary isoclinic irregular type at 00 and arbitrary regular monodromy at 00 (when $\nu\geq 1$, the monodromy can be arbitrary).
  • (ii)
    If we regard the equation in (43) as with coefficients in $\mathbb{C}[t,t^{-1}]$, then they will have regular singularity at $\infty$, but with not necessarily trivial monodromy.
6. Central support of toral K-types
The counterpart on the automorphic side of irreducible isoclinic formal connections will be the so-called toral supercuspidal representations [KalethaRegular, §4.8]. This family of supercuspidal representations are originally due to Adler [Adler], and contain epipelagic representations as special cases. We give a brief review of them following the exposition in [FintzenIHES, §3.6].
6.1. Recollection on toral supercuspidal representations
In one word, toral supercuspidal representations are those supercuspidal representations induced from Yu data of the form $(S\subset G;x;\nu;\phi,1)$, where the only proper twisted Levi subgroup is an elliptic torus. To make it more explicit, we carry out the construction in §3 in this situation.
6.1.1.
Let $m$ be a regular elliptic number of $W=N_{G}(T)/T$. Let $\check{\rho}_{G}\in X_{*}(T/Z(G))$ be the half sum of positive coroots, and $x$ the point in the closure of fundamental alcove of Bruhat-Tits building $\mathcal{B}(G,T)$ that is conjugated to $\frac{\check{\rho}_{G}}{m}$. Write $\lambda=mx\in X_{*}(T/Z(G))$, which is conjugated to $\check{\rho}_{G}$. Then $\Theta=\mathrm{Ad}_{\lambda(\zeta_{m})}$ defines a grading
$\mathfrak{g}=\bigoplus_{i\in\mathbb{Z}/m\mathbb{Z}}\mathfrak{g}_{i}.$
Let $\nu=\frac{N}{m}$ where $(N,m)=1$. As on the spectral side, there exists regular semisimple element $Y\in\mathfrak{g}_{-N}^{\mathrm{rs}}\simeq\mathfrak{g}_{N}^{*,\mathrm{rs}}$. Denote $T_{Y}=C_{G}(Y)$, $\mathfrak{t}_{Y}=\mathfrak{z}_{\mathfrak{g}}(Y)$. Since $Y$ is $\Theta$-eigen, $\Theta$ preserves $\mathfrak{t}_{Y}$ and induces grading
$\mathfrak{t}_{Y}=\bigoplus_{i\in\mathbb{Z}/m\mathbb{Z}}\mathfrak{t}_{Y,i}.$
Note that since the action of $\Theta$ on $T_{Y}$ is a regular elliptic element of order $m$ in the Weyl group $W(G,T_{Y})$, we have $\mathfrak{t}_{Y,0}=0$.
Fix $u=t^{1/m}$ and denote $E=\mathbb{C}(\!(u)\!)$. We have a Moy-Prasad grading on $\mathfrak{g}(\!(t)\!)$:
(46) $\psi:\mathfrak{g}(\!(t)\!)\simeq\widehat{\bigoplus_{i\in\mathbb{Z}}}\mathfrak{g}_{i}u^{i},\quad X(t)\mapsto\mathrm{Ad}_{\lambda(u)}X(u^{m}).$
Denote $r_{m}:u\mapsto\zeta_{m}^{-1}u$ and $\sigma=\Theta\times r_{m}\in\mathrm{Aut}(\mathfrak{g}(E))$. The RHS of the above is $\mathfrak{g}(E)^{\sigma}$. Henceforth we will omit $\psi$ and identify $\mathfrak{g}(\!(t)\!)$ with ^$\widehat{\bigoplus}_{i\in\mathbb{Z}}\mathfrak{g}_{i}u^{i}$. Denote $\mathfrak{g}(\!(t)\!)_{i}\simeq\mathfrak{g}_{i}u^{i}$ and $\mathfrak{g}(\!(t)\!)_{\geq i}\simeq\widehat{\bigoplus}_{j\geq i}\mathfrak{g}(\!(t)\!)_{i}$.
The element $Y\in\mathfrak{g}_{-N}^{\mathrm{rs}}$ defines an element $Y^{\prime}=\psi^{-1}(Yu^{-N})$, which is regular semisimple in $\mathfrak{g}(\!(t)\!)$. Let $S=C_{G(\!(t)\!)}(Y^{\prime})$ be the centralizer and $\mathfrak{s}=\mathfrak{z}_{\mathfrak{g}(\!(t)\!)}(Y^{\prime})$. Since $S\simeq(\mathrm{Res}_{E/F}T_{Y}\times_{\mathbb{C}}E)^{\Theta\times r_{m}^{-1}}$, $S$ is an elliptic torus. We have
$\mathfrak{s}\simeq\widehat{\bigoplus_{i\in\mathbb{Z}}}\mathfrak{t}_{Y,i}u^{i}.$
Note that $x=\lambda/m$ is the unique point in the Bruhat-Tits building of $S$. Denote by $S_{r}$ the Moy-Prasad subgroup of $S$ with depth $r\in\mathbb{R}_{\geq 0}$ and $S_{r+}=\bigcup_{s>r}S_{s}$.
Now take arbitrary elements $Y_{i}\in\mathfrak{t}_{Y,i}$ for $-N+1\leq i\leq-1$ and denote $Y_{-N}=Y$. Then $\psi^{-1}(\sum_{i=-N}^{-1}Y_{i}u^{i})$ determines a character of $S_{0+}$ of depth $\nu$. Let $\phi:S\rightarrow\mathbb{C}^{\times}$ be a character whose differential is as above. Such $(S,\phi)$ gives a tame elliptic regular pair of depth $\nu$ [KalethaRegular, Definition 3.7.5].
6.1.2.
We need to extend $\phi$ further. Let $P=G_{x}$ be the parahoric subgroup determined by $x$, and $P(i)=G_{x,i/m}$ its Moy-Prasad subgroups. Denote their Lie algebras by $\mathfrak{p},\mathfrak{p}(i)$. Denote $S(i)=S\cap P(i)$, $\mathfrak{s}(i)=\mathfrak{s}\cap\mathfrak{p}(i)$. Note $S(1)=S_{0+}$. The construction varies slightly depending on the parity of $N$.
Case 1: $N$ is odd.
In this case $G_{x,\frac{\nu}{2}}=G_{x,\frac{\nu}{2}+}=P(\frac{N+1}{2})$. Consider the root subspaces with respect to $\mathfrak{t}_{Y}$:
(47) $\mathfrak{g}=\mathfrak{t}_{Y}\oplus\tau,\qquad\tau=\bigoplus_{\alpha\in\Phi(\mathfrak{g},\mathfrak{t}_{Y})}\mathfrak{g}_{\alpha}.$
The automorphism acts on $\mathfrak{t}_{Y}$ and $\Phi(\mathfrak{g},\mathfrak{t}_{Y})$, thus acts on $\tau$. Let $\tau_{i}=\tau\cap\mathfrak{g}_{i}$, then
$\mathfrak{g}_{i}=\mathfrak{t}_{Y,i}\oplus\tau_{i}.$
Here although the notations $\mathfrak{g}_{i},\tau_{i}$ coincide with the Lie algebras of twisted Levi subalgebras in a general Yu datum, this should raise no confusion.
Since $P(\frac{N+1}{2})/P(N+1)$ is abelian and unipotent, $P(\frac{N+1}{2})/P(N+1)\simeq\mathfrak{p}(\frac{N+1}{2})/\mathfrak{p}(N+1)$. Consider projection
$\pi:P(\frac{N+1}{2})\twoheadrightarrow$ $P(\frac{N+1}{2})/P(N+1)\simeq\mathfrak{p}(\frac{N+1}{2})/\mathfrak{p}(N+1)\simeq\bigoplus_{i=\frac{N+1}{2}}^{N}\mathfrak{g}_{i}u^{i}$
[…]
Let $J=S(1)P(\frac{N+1}{2})$. Define a character $\tilde{\phi}$ by $\tilde{\phi}(G_{x,\nu+})=1$, $\tilde{\phi}|_{S(1)}=\phi$, $\tilde{\phi}|_{G_{x,\frac{\nu}{2}}}=\phi\circ\pi$.
Case 2: $N$ is even.
Note that $Y\in\mathfrak{g}_{N}^{*}$ induces a non-degenerate symplectic form on $G_{x,\frac{\nu}{2}}/S(\frac{N}{2})G_{x,\frac{\nu}{2}}+\simeq\mathfrak{g}_{\frac{N}{2}}/\mathfrak{s}_{\frac{N}{2}}\simeq\tau_{\frac{N}{2}}$ by
(48) $(x,y)\in\tau_{\frac{N}{2}}\times\tau_{\frac{N}{2}}\rightarrow\langle Y,[x,y]\rangle.$
We choose a Lagrangian subspace $\mathfrak{m}\subset\tau_{\frac{N}{2}}$ as follows. Recall that the order $m$ automorphism $\Theta$ is regular elliptic, so that the subgroup of automorphisms generated by $\Theta$ acts freely on $\Phi(\mathfrak{g},\mathfrak{t}_{Y})$. For each $\Theta$-orbit $O\subset\Phi(\mathfrak{g},\mathfrak{t}_{Y})$, pick a representative $\alpha_{O}\in O$. Fix a basis $X_{\alpha_{O}}\in\mathfrak{g}_{\alpha_{O}}$. Then a basis of $\tau_{\frac{N}{2}}$ can be given by
(49) $X_{O,\frac{N}{2}}=X_{\alpha_{O}}+\zeta_{m}^{-\frac{N}{2}}\Theta(X_{\alpha_{O}})+\zeta_{m}^{-N}\Theta^{2}(X_{\alpha_{O}})+\cdots+\zeta_{m}^{-(m-1)\frac{N}{2}}\Theta^{m-1}(X_{\alpha_{O}}).$
Here $\Theta^{j}(X_{\alpha_{O}})\in\mathfrak{g}_{\Theta^{j}(\alpha_{O})}$. We claim that for any root $\alpha$ and $j$, $\Theta^{j}\alpha\neq-\alpha$. Otherwise $\Theta^{2j}\alpha=\alpha$, so that $m|2j$. Since $N$ is even, $(N,m)=1$, thus $m$ is odd, $m|j$, $\Theta^{j}\alpha=\alpha\neq-\alpha$.
Now we define a partition $\Theta\backslash\Phi(\mathfrak{g},\mathfrak{t}_{Y})=A\sqcup B$ of orbits as follows. Take any orbit $O$ and put it in $A$. For any $\alpha\in O$, from previous discussion we see $-\alpha\not\in O$, and we put the orbit of $-\alpha$ in B. Repeat this for the rest of orbits. The resulting partition satisfies $|A|=|B|=\frac{1}{2}\dim\tau_{\frac{N}{2}}$, and any $\alpha,\beta$ in the union of orbits in $A$ satisfy $\alpha\neq-\beta$. We let
(50) $\mathfrak{m}=\bigoplus_{O\in A}\mathbb{C}X_{O,\frac{N}{2}}.$
Thus $\dim\mathfrak{m}=\frac{1}{2}\dim\tau_{\frac{N}{2}}$. For any $O,O^{\prime}\in A$, since $\alpha\neq-\beta$ for any $\alpha\in O$ and $\beta\in O^{\prime}$, we have $[X_{O,\frac{N}{2}},X_{O^{\prime},\frac{N}{2}}]\in\tau\cap\mathfrak{g}_{N}=\tau_{N}$. Since the pairing of $\mathfrak{t}_{Y}$ with $\tau$ vanishes, $\mathfrak{m}$ is isotropic with respect to (48), thus a Lagrangian of $\tau_{\frac{N}{2}}$.
Let $P(\frac{N}{2})_{\mathfrak{m}}$ be the preimage of $\mathfrak{m}$ under $P(\frac{N}{2})/S(\frac{N}{2})P(\frac{N}{2}+1)\simeq\tau_{\frac{N}{2}}$. Let $J=S(1)P(\frac{N}{2})_{\mathfrak{m}}$ and let $\tilde{\phi}:J\rightarrow\mathbb{C}^{\times}$ be the composition of $\phi$ with projection to $S(1)$.
6.1.3.
We call $(J,\tilde{\phi})$ a toral supercuspidal K-type. It is the level structure associated to Yu datum $(S\subset G;x;\nu;\phi,1)$. If we replace loop group with $p$-adic group and the choice of Lagrangian with Heisenberg and Weil representation, then as a special case of [FintzenTame, Theorem 3.1], the induced representation $\mathrm{c-ind}^{G(F)}_{J}\tilde{\phi}$ is irreducible, hence supercuspidal.
6.2. Segal-Sugawara operators
Recall $\widehat{\mathfrak{g}}=\mathfrak{g}(\!(t)\!)\oplus\mathbb{C}\cdot\mathbf{1}$ is the affine Kac-Moody algebra at the critical level associated to $\mathfrak{g}$. The element $\mathbf{1}$ is central, and the Lie bracket is given by
(51) $[A\otimes f(t),B\otimes g(t)]=[A,B]\otimes f(t)g(t)-\frac{1}{2}\mathrm{tr}(\mathrm{ad}_{A}\circ\mathrm{ad}_{B})\mathrm{Res}(gdf)\cdot\mathbf{1}.$
Recall that $\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})$ has coordinates $v_{i,j}$, $1\leq i\leq n,j\in\mathbb{Z}$. Let $S_{i,j}\in\mathfrak{Z}$ be the image of $v_{ij}$ via the above isomorphism. This gives a set of topological generators of $\mathfrak{Z}$ called Segal-Sugawara operators, c.f. [FrenkelBook, §4.3.1., §4.3.2.].
To write down $S_{i,j}$ more explicitly, we fix a topological basis of $\widehat{U}(\mathfrak{g})$ as follows. First we fix a topological basis of $\mathfrak{g}(\!(t)\!)$. Recall the (completed) grading given by (46). Each $\mathfrak{g}(\!(t)\!)_{i}\simeq\mathfrak{g}_{i}u^{i}=\mathfrak{t}_{Y,i}u^{i}\oplus\tau_{i}u^{i}$. Fix arbitrary basis $A_{i,k}$, $1\leq k\leq\dim\mathfrak{t}_{Y,i}$ of $\mathfrak{t}_{Y,i}$. For $\tau_{i}$, for each $\Theta$-orbit $O$ on $\Phi(\mathfrak{g},\mathfrak{t}_{Y})$, pick a representative $\alpha_{O}\in O$. Fix a basis $X_{\alpha_{O}}\in\mathfrak{g}_{\alpha_{O}}$. Similarly as for $i=\frac{N}{2}$, a basis of $\tau_{i}$ can be given by
(52) $X_{O,i}=X_{\alpha_{O}}+\zeta_{m}^{-i}\Theta(X_{\alpha_{O}})+\zeta_{m}^{-2i}\Theta^{2}(X_{\alpha_{O}})+\cdots+\zeta_{m}^{-(m-1)i}\Theta^{m-1}(X_{\alpha_{O}}).$
In particular, $\dim\tau_{i}=|\Theta\backslash\Phi(\mathfrak{g},\mathfrak{t}_{Y})|=\frac{|\Phi|}{m}$ is constant as seen in [Panyushev, Theorem 4.2.(ii)]. Moreover, :=$\prod:=\{A_{i,k}u^{i},X_{O,i}u^{i}\mid i\in\mathbb{Z},1\leq k\leq\dim\mathfrak{t}_{Y,i},O\in\Theta\backslash\Phi(\mathfrak{g},\mathfrak{t}_{Y})\}$ gives a basis of $\mathfrak{g}(\!(t)\!)$. We put a total order on $\prod$ by requiring
$\cdots<\mathfrak{g}(\!(t)\!)_{i-1}<\mathfrak{g}(\!(t)\!)_{i}<\mathfrak{g}(\!(t)\!)_{i+1}<\cdots,$
$\tau_{Y,i}u^{i}<\mathfrak{t}_{Y,i}u^{i}$, and fix arbitrary total order on $\{A_{i,k}u^{i}\}_{k}$ and $\{X_{O,i}u^{i}\}_{O}$. By PBW theorem, a topological basis of $\widehat{U}(\mathfrak{g})$ can be given by
(53) $\{X_{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{k}}\mid\ k\geq 0,\ X_{a_{i}}\in\prod,\ X_{a_{1}}\leq X_{a_{2}}\leq\cdots\leq X_{a_{k}}\}.$
Lemma 27.
The operator $S_{i,j}\in\mathfrak{Z}$ acts on a vector in a smooth $\widehat{\mathfrak{g}}$-module as a finite sum of tensors
$X_{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{k}}$
satisfying
$\mathrm{(i)}\ X_{a_{r}}\in\mathfrak{g}_{a_{r}}u^{a_{r}}\cap\prod,\qquad\mathrm{(ii)}\ k\leq d_{i},\qquad\mathrm{(iii)}\ \sum_{r=1}^{k}a_{r}=m(j-d_{i}+1).$
Proof.
As in [KXY, Lemma 54], $S_{i,j}$ acts as a finite linear combination of tensors
$E_{\beta_{1}}t^{n_{1}}\otimes\cdots\otimes E_{\beta_{k}}t^{n_{k}}$
where $k\leq d_{i}$, $E_{\beta_{r}}$ is a basis of $\mathfrak{g}_{\beta_{r}}$ with respect to $\mathfrak{t}$, $\sum_{r=1}^{k}n_{r}=j-d_{i}+1$, $\sum_{r=1}^{k}\beta_{r}=0$. Rewriting each $E_{\beta_{1}}t^{n_{1}}$ in terms of basis $\prod$, we obtain a linear combination of tensors satisfying (i), (ii) of the lemma. Moreover, note that the isomorphism $\psi$ is by replacing $t=u^{m}$ and applying $\mathrm{Ad}_{\lambda(u)}$. We obtain
$\sum_{r=1}^{k}a_{r}=m\sum_{r=1}^{k}n_{r}+\sum_{r=1}^{k}\langle\lambda,\beta_{r}\rangle=m(j-d_{i}+1).$
6.3. Central characters
Recall we defined a subspace $\mathfrak{j}^{\prime}\subset\mathfrak{j}$ (30). Explicitly, in the current situation it is given by
$\mathfrak{j}^{\prime}=\begin{cases}\bigoplus_{i=\frac{N+1}{2}}^{N}\tau_{i}u^{i}\oplus\mathfrak{p}(N+1),\hskip 49.79231ptN\text{ odd,}\\ \bigoplus_{i=\frac{N}{2}+1}^{N}\tau_{i}u^{i}\oplus\mathfrak{m}u^{\frac{N}{2}}\oplus\mathfrak{p}(N+1),\quad N\text{ even.}\end{cases}$
Lemma 28.
The subspace $\mathfrak{j}^{\prime}\subset\mathfrak{j}$ is an ideal, so that $\mathfrak{j}^{\prime}=\mathfrak{j}^{+}$.
Proof.
We need to show $[\mathfrak{j},\mathfrak{j}^{\prime}]\subset\mathfrak{j}^{\prime}$. Observe $\mathfrak{j}=\mathfrak{j}^{\prime}+\mathfrak{s}(1)$. It is easy to see that the adjoint action of $\mathfrak{s}(1)$ preserves $\mathfrak{j}^{\prime}$. To show $[\mathfrak{j}^{\prime},\mathfrak{j}^{\prime}]\subset\mathfrak{j}^{\prime}$, from the depth consideration it suffices to show $[\mathfrak{m},\mathfrak{m}]\subset\tau_{N}$ when $N$ is even. This follows from the construction of $\mathfrak{m}$ (50). ∎
Therefore, the quotient (31) becomes $\bar{\mathfrak{j}}=\mathfrak{j}/\mathfrak{j}^{+}\simeq\mathfrak{s}(1)/\mathfrak{s}(N+1)\simeq\bigoplus_{1\leq i\leq N}\mathfrak{t}_{Y,i}u^{i}$.
As in §4.2.2, $\mathfrak{j}^{+}\subset\mathfrak{j}\subset\mathfrak{g}[\![t]\!]$, $\mathfrak{j}^{+}+\mathbb{C}\mathbf{1}\subset\mathfrak{j}+\mathbb{C}\mathbf{1}$ are subalgebras of $\widehat{\mathfrak{g}}$. Consider the $\widehat{\mathfrak{g}}$-module $\mathrm{Vac}_{\mathfrak{j}^{+}}$ defined in (32).
We have
$\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}}\simeq\mathrm{Hom}_{\mathfrak{j}}(U(\bar{\mathfrak{j}}),\widehat{U}(\mathfrak{g})\otimes_{U(\mathfrak{j}+\mathbb{C}\mathbf{1})}U(\bar{\mathfrak{j}}))\simeq\mathrm{Vac}_{\mathfrak{j}^{+}}^{J^{+}}.$
Since $\bar{\mathfrak{j}}\simeq\mathfrak{s}(1)/\mathfrak{s}(N+1)$ is abelian, $U(\bar{\mathfrak{j}})\hookrightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}}$. Recall $\mathfrak{Z}_{\mathfrak{j}^{+}}=\mathrm{Im}(\mathfrak{Z}\rightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}})$. Define
(54) $A^{\prime}:=\mathbb{C}[S_{i,j}\mid(i,j)\in\tilde{A}_{\nu}]$
where $\tilde{A}_{\nu}$ is as in (41).
Proposition 29.
The image of $A^{\prime}\to\mathfrak{Z}_{\mathfrak{j}^{+}}$ is contained in $U(\bar{\mathfrak{j}})$, so that we have commutative diagram
[…]
Proof.
The action of central operators $S_{i,j}$ on $\mathrm{Vac}_{\mathfrak{j}^{+}}$ is determined by their actions on $1\in\mathrm{Vac}_{\mathfrak{j}^{+}}$. By Lemma 27, $S_{i,j}\cdot 1$ can be written as a finite linear combination of tensors
$X_{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{k}},$
where $k\leq d_{i}$, $X_{a_{r}}\in\mathfrak{g}_{a_{r}}u^{a_{r}}\cap\prod$, $\sum_{r=1}^{k}a_{r}=m(j-d_{i}+1)$, $a_{1}\leq a_{2}\leq\cdots\leq a_{k}$.
For $(i,j)\in\tilde{A}_{\nu}$,
$(d_{i}-1)N+1\leq\sum_{r=1}^{k}a_{r}=m(j-d_{i}+1)=(d_{i}-1)N-\ell_{i,j}\leq d_{i}N.$
Note that for $a\geq N+1$, $\mathfrak{g}_{a}u^{a}\subset\mathfrak{p}(N+1)\subset\mathfrak{j}^{+}$ annihilates $1\in\mathrm{Vac}_{\mathfrak{j}^{+}}$. Thus if $a_{k}\geq N+1$, the tensor has zero coefficient in $S_{i,j}\cdot 1$. We can assume $a_{1}\leq\cdots\leq a_{k}\leq N$. Then
$a_{1}\geq\sum_{r=1}^{k}a_{r}-(k-1)N\geq(d_{i}-1)N+1-(d_{i}-1)N=1.$
Suppose $k\leq d_{i}-1$. Then
$a_{1}\geq\sum_{r=1}^{k}a_{r}-(k-1)N\geq(d_{i}-1)N+1-(d_{i}-2)N=N+1,$
contradicting with assumption $a_{1}\leq a_{k}\leq N$. So $k=d_{i}$.
Assume
$1\leq a_{1}\leq a_{2}\leq\cdots\leq a_{p}\leq\frac{N}{2}<\frac{N+1}{2}\leq a_{p+1}\leq\cdots\leq a_{d_{i}}.$
If $p=0$, then all $a_{r}>\frac{N}{2}$, $X_{a_{r}}\in\mathfrak{g}_{a_{r}}u^{a_{r}}\subset\mathfrak{j}^{+}$. The tensor is contained in $U(\bar{\mathfrak{j}})$.
If $p>0$, we have inequality
$(d_{i}-1)N+1\leq\sum_{r=1}^{d_{i}}a_{r}\leq p\lfloor\frac{N}{2}\rfloor+(d_{i}-p)N.$
We get
$p\leq\frac{N-1}{N-\lfloor\frac{N}{2}\rfloor}=\begin{cases}2\frac{N-1}{N+1},\quad N\text{ odd}\\ 2\frac{N-1}{N},\quad N\text{ even}\end{cases}<2.$
We obtain $p=1$ and $a_{2},...,a_{d_{i}}\geq\frac{N+1}{2}$, $X_{a_{r}}\in\mathfrak{j}$ for $r\geq 2$.
If $X_{a_{d_{i}}}\in\tau_{a_{d_{i}}}u^{a_{d_{i}}}$, then the tensor acts as 00 on $1$. Thus we can assume $X_{a_{d_{i}}}\in\mathfrak{t}_{Y,a_{d_{i}}}u^{a_{d_{i}}}$. We show $X_{a_{r}}\in\mathfrak{t}_{Y,a_{r}}u^{a_{r}}$ for all $2\leq r\leq d_{i}$ by induction. Given $2\leq q<d_{i}$, suppose $X_{a_{r}}\in\mathfrak{t}_{Y,a_{r}}u^{a_{r}}$ for all $q<r\leq d_{i}$. If $X_{a_{q}}\in\tau_{a_{q}}u^{a_{q}}$, we claim the tensor acts as 00 on $1$. In fact, we can move $X_{a_{q}}$ to the right most term, after which the tensor acts as 00 on $1$. In this process, two kinds of shorter tensors are produced by (51): replacing two adjacent terms in the tensor either by a multiple of unit $\mathbf{1}$, which gives a length $d_{i}-2$ tensor, or by the Lie bracket of the two terms, which gives a length $d_{i}-1$ tensor. However, since the Killing form between $\mathfrak{t}_{Y}$ and $\tau$ vanishes, only the second case can contribute nonzero tensors, where the sum $\sum_{r=1}^{k}a_{r}$ is the same as that of the original tensor. We can repeat this process to rewrite these shorter tensors as a linear combination of basis in $\prod$. From the procedure, these tensors all satisfy the conditions on $X_{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{k}}$, and moreover $k<d_{i}$. From previous discussion, we know such shorter tensors all act as 00 on $1$. We conclude that $X_{a_{r}}\in\mathfrak{t}_{Y,a_{r}}u^{a_{r}}$ for all $2\leq r\leq d_{i}$.
It remains to show $X_{a_{1}}\in\mathfrak{t}_{Y,a_{1}}u^{a_{1}}$. Suppose $X_{a_{1}}\in\tau_{a_{1}}u^{a_{1}}$ instead. Note that the central element $S_{i,j}$ commutes with (the preimage in $\widehat{U}(\mathfrak{g})$ of) $\mathfrak{t}_{Y,d}u^{d}$ for all $d$. Let $\widehat{U}(\mathfrak{g})_{\leq e}$ be the subspace of tensors in $\widehat{U}(\mathfrak{g})$ with length $\leq e$, and $\mathrm{gr}_{e}\widehat{U}(\mathfrak{g}):=\widehat{U}(\mathfrak{g})_{\leq e}/\widehat{U}(\mathfrak{g})_{\leq e+1}$. Note that $\mathrm{gr}_{e}\widehat{U}(\mathfrak{g})$ has a basis by images of tensors as in Lemma 27 with $k=e$. Denote by $\mathrm{gr}_{e}^{d}\widehat{U}(\mathfrak{g})$ the subspace of $\mathrm{gr}_{e}\widehat{U}(\mathfrak{g})$ generated by tensors satisfying $\sum_{r=1}^{k}a_{r}=d$, and denote its basis as above by $C(e,d)$. The action of $\mathfrak{t}_{Y,d}u^{d}$ sends $\mathrm{gr}_{d_{i}}^{m(j-d_{i}+1)}\widehat{U}(\mathfrak{g})$ to $\mathrm{gr}_{d_{i}}^{m(j-d_{i}+1+d)}\widehat{U}(\mathfrak{g})$. Moreover, this action sends the subspace of $\mathrm{gr}_{d_{i}}^{m(j-d_{i}+1)}\widehat{U}(\mathfrak{g})$ spanned by
$\tau_{a_{1}}u^{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{d_{i}}}$
into the subspace spanned by
$\tau_{a_{1}+d}u^{a_{1}+d}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{d_{i}}}.$
It also sends those basis in $C(d_{i},m(j-d_{i}+1))$ not contained in $\tau_{a_{1}}u^{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{d_{i}}}$ into the subspace generated by those basis in $C(d_{i},m(j-d_{i}+1+d))$ not contained in $\tau_{a_{1}+d}u^{a_{1}+d}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{d_{i}}}$.
Let $Zu^{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{d_{i}}}$, $Z\in\tau_{a_{1}}$ be the component of the image of $S_{i,j}$ in $\mathrm{gr}_{d_{i}}^{m(j-d_{i}+1)}\widehat{U}(\mathfrak{g})$ inside $\tau_{a_{1}}u^{a_{1}}\otimes X_{a_{2}}\otimes\cdots\otimes X_{a_{d_{i}}}$ in terms of the fixed basis. From the above discussion together with $[\mathfrak{t}_{Y,d}u^{d},S_{i,j}]=0$, we obtain
$[\mathfrak{t}_{Y,d}u^{d},Zu^{a_{1}}]=0,\quad\forall d.$
Thus $[\mathfrak{t}_{Y,d},Z]=0,\forall d$, so that $[\mathfrak{t}_{Y},Z]=0$. Since $Z\in\tau$, we conclude $Z=0$. This completes the proof. ∎
6.4. Local Hitchin image at level $\mathfrak{j}^{+}$
We prepare some results on local Hitchin images demanded for a more detailed study of (55).
6.4.1.
We first fix a Kostant section of $\mathfrak{g}$ as follows. Recall in §6.2 we have fixed a principal $\mathfrak{sl}_{2}$-triple $\{p_{-1},2\check{\rho}_{\check{G}},p_{1}\}$ in $\check{\mathfrak{g}}$ together with a basis of Kostant section $p_{-1}+\check{\mathfrak{g}}^{p_{1}}=p_{-1}+\sum_{i=1}^{n}\mathbb{C}p_{i}$. Consider the dual principal $\mathfrak{sl}_{2}$-triple $\{p_{-1},2\check{\rho}_{G},p_{1}\}$ in $\mathfrak{g}$ where we abuse of notations. We have isomorphisms
$p_{-1}+\check{\mathfrak{g}}^{p_{1}}\simeq\check{\mathfrak{g}}/\!\!/\check{G}\simeq\check{\mathfrak{t}}/\!\!/W\simeq\mathfrak{t}^{*}/\!\!/W\simeq\mathfrak{t}/\!\!/W\simeq\mathfrak{g}/\!\!/G\simeq p_{-1}+\mathfrak{g}^{p_{1}}$
where $\mathfrak{t}\simeq\mathfrak{t}^{*}$ is induced from Killing form. The above morphisms are compatible with $\mathbb{G}_{m}$-action, where on $p_{-1}+\check{\mathfrak{g}}^{p_{1}}$ the action is given by $\check{\rho}_{\check{G}}$ on $\check{\mathfrak{g}}^{p_{1}}$ and on $p_{-1}+\mathfrak{g}^{p_{1}}$ the action is given by $\check{\rho}_{G}$ on $\mathfrak{g}^{p_{1}}$. The weight subspaces of $\check{\mathfrak{g}}^{p_{1}}$ are isomorphic to those of $\mathfrak{g}^{p_{1}}$, and the homogeneous basis $p_{i}$ of $\check{\mathfrak{g}}^{p_{1}}$ are sent to a homogeneous basis of $\mathfrak{g}^{p_{1}}$ which we still denote by $p_{i}$.
Consider the local Hitchin map
(56) $h^{cl}:\mathfrak{g}^{*}(\!(t)\!)\mathrm{d}t\simeq\mathfrak{g}(\!(t)\!)\mathrm{d}t\rightarrow\mathrm{Hit}(D^{\times})\simeq\bigoplus_{i=1}^{n}\mathbb{C}(\!(t)\!)(\mathrm{d}t)^{d_{i}},$
where the last isomorphism is via invariant polynomials defined by Kostant section $p_{-1}+\sum_{i}\mathbb{C}p_{i}\simeq\mathfrak{g}/\!/G$. Let $h_{i,j}$ be the coordinate of $\mathbb{C}t^{-j-1}(\mathrm{d}t)^{d_{i}}$. Under the isomorphism $\mathrm{gr}\mathfrak{Z}\simeq\mathrm{Fun}~{}\mathrm{Hit}(D^{\times})$, the image of $S_{i,j}$ is $h_{i,j}$.
6.4.2.
Let $\mathfrak{j}^{+,\perp}\subset\mathfrak{g}^{*}(\!(t)\!)\mathrm{d}t$ be the $\mathcal{O}$-lattice that is orthogonal to $\mathfrak{j}^{+}$, i.e.
$\mathfrak{j}^{+,\perp}=\{X\mathrm{d}t\in\mathfrak{g}^{*}(\!(t)\!)\mathrm{d}t\mid\mathrm{Res}\langle X,Y\rangle\mathrm{d}t=0,\ \forall\ Y\in\mathfrak{j}^{+}\}.$
With respect to the filtration induced from $\widehat{U}(\mathfrak{g})$, $\mathrm{gr}\mathrm{Vac}_{\mathfrak{j}^{+}}\simeq\mathrm{Fun}~{}\mathfrak{j}^{+,\perp}$. Identifying $\mathfrak{g}^{*}\simeq\mathfrak{g}$ using Killing form and applying $\psi$, we have
(57) $\mathfrak{j}^{+,\perp}\simeq\begin{cases}(\bigoplus_{-N\leq i\leq-\frac{N+1}{2}}\mathfrak{t}_{Y,i}u^{i}\oplus\prod_{i\geq-\frac{N-1}{2}}\mathfrak{g}_{i}u^{i})\frac{\mathrm{d}u}{u},\hskip 76.82234ptN\text{ odd,}\\[5.69046pt] (\bigoplus_{-N\leq i\leq-\frac{N}{2}-1}\mathfrak{t}_{Y,i}u^{i}\oplus\mathfrak{m}^{\perp}u^{-\frac{N}{2}}\oplus\prod_{i\geq-\frac{N}{2}+1}\mathfrak{g}_{i}u^{i})\frac{\mathrm{d}u}{u},\qquad N\text{ even,}\end{cases}$
where $\mathfrak{t}_{Y,-\frac{N}{2}}\subset\mathfrak{m}^{\perp}\subset\mathfrak{g}_{-\frac{N}{2}}$ is the orthogonal complement of $\mathfrak{m}$ with respect to Killing form.
Similarly define lattice $\mathfrak{j}^{\perp}$ with isomorphism
(58) $\mathfrak{j}^{\perp}\simeq\begin{cases}(\bigoplus_{-\frac{N-1}{2}\leq i\leq-1}\tau_{i}u^{i}\oplus\prod_{i\geq 0}\mathfrak{g}_{i}u^{i})\frac{\mathrm{d}u}{u},\hskip 116.65646ptN\text{ odd,}\\[5.69046pt] ((\tau_{-\frac{N}{2}}\cap\mathfrak{m}^{\perp})u^{-\frac{N}{2}}\oplus\bigoplus_{-\frac{N}{2}+1\leq i\leq-1}\tau_{i}u^{i}\oplus\prod_{i\geq 0}\mathfrak{g}_{i}u^{i})\frac{\mathrm{d}u}{u},\qquad N\text{ even.}\end{cases}$
Note that $\mathfrak{j}^{+,\perp}/\mathfrak{j}^{\perp}\simeq(\bigoplus_{-N\leq i\leq-1}\mathfrak{t}_{Y,i}u^{i})\frac{\mathrm{d}u}{u}\simeq\bar{\mathfrak{j}}^{*}$.
Denote by $\mathrm{Hit}(D)_{\mathfrak{j}^{+}}$ the closure of the image $h^{cl}(\mathfrak{j}^{+,\perp})$, then $\mathrm{gr}\mathfrak{Z}_{\mathfrak{j}^{+}}\simeq\mathrm{Fun}~{}\mathrm{Hit}(D)_{\mathfrak{j}^{+}}$.
Define
(59) $\mathrm{Hit}(D)_{\mathfrak{j}^{+}}^{\prime}:=\bigoplus_{i=1}^{n}t^{-d_{i}-\lfloor\frac{d_{i}N}{m}\rfloor)}\mathbb{C}[\![t]\!](\mathrm{d}t)^{d_{i}}.$
Proposition 30.
$\mathrm{Hit}(D)_{\mathfrak{j}^{+}}=\mathrm{Hit}(D)_{\mathfrak{j}^{+}}^{\prime}$.
Proof.
Since $\mathfrak{j}^{+}\supset\mathfrak{p}(N+1)$, $\mathfrak{j}^{+,\perp}\subset\mathfrak{p}(N+1)^{\perp}$. By [Zhu, Proposition 10],
(60) $h^{cl}(\mathfrak{p}(N+1)^{\perp})\subset\mathrm{Hit}(D)_{\mathfrak{j}^{+}}^{\prime}.$
It remains to show the closure of $h^{cl}(\mathfrak{j}^{+,\perp})$ contains $\mathrm{Hit}(D)_{\mathfrak{j}^{+}}^{\prime}$.
Since $\Theta$ is $\mathbb{Z}$-regular elliptic [RLYG, §2], we know by [RLYG, Proposition 8] that $\mathrm{Ad}_{\check{\rho}(\zeta_{m})}$ and $\mathrm{Ad}_{\check{\rho}(\zeta_{m}^{N})}$ are conjugated, so are $\Theta=\mathrm{Ad}_{\lambda(\zeta_{m})}$ are $\Theta^{N}=\mathrm{Ad}_{\lambda(\zeta_{m}^{N})}$. Thus we can find a $\Theta$-adapted principal $\mathfrak{sl}_{2}$-triple $\{e,h,f\}$ with $e\in\mathfrak{g}_{N},f\in\mathfrak{g}_{-N},h\in\mathfrak{g}_{0}$. Let $g_{0}\in G$ be the unique element up to center that conjugates $\{p_{1},2\check{\rho},p_{-1}\}$ to $\{e,h,f\}$. Denote $p_{i}^{\prime}=g_{0}\cdot p_{i}\in\mathfrak{g}_{(d_{i}-1)N}$.
An element $\sum_{i=1}^{n}\sum_{j\leq d_{i}-1+\lfloor\frac{d_{i}N}{m}\rfloor}h_{i,j}t^{-j-1}(\mathrm{d}t)^{d_{i}}=\sum_{i=1}^{n}t^{-\lfloor\frac{d_{i}N}{m}\rfloor}c_{i}(t)(\frac{\mathrm{d}t}{t})^{d_{i}}\in\mathrm{Hit}(D)_{\mathfrak{j}^{+}}^{\prime}$, rewritten in terms of $u$, is the image of
$(p_{-1}+\sum_{i=1}^{n}u^{-d_{i}-m\lfloor\frac{d_{i}N}{m}\rfloor}c_{i}(u^{m})p_{i})\mathrm{d}u.$
If $N$ is odd, applying conjugation by $g_{0}$, we get
$(f+\sum_{i=1}^{n}u^{-d_{i}-m\lfloor\frac{d_{i}N}{m}\rfloor}c_{i}(u^{m})p_{i}^{\prime})\mathrm{d}u.$
Applying conjugation by $h(u^{\frac{N+1}{2}})$, we get
(61) $mu^{-N}(f+\sum_{i=1}^{n}u^{d_{i}N-m\lfloor\frac{d_{i}N}{m}\rfloor}c_{i}(u^{m})p_{i}^{\prime})\frac{\mathrm{d}u}{u}\in\prod_{i\geq-N}\mathfrak{g}_{i}u^{i}\frac{\mathrm{d}u}{u},$
where $c_{i}(t)\in\mathbb{C}[\![t]\!]$ with $c_{i}(0)=h_{i,d_{i}-1+\lfloor\frac{d_{i}N}{m}\rfloor}$.
If $N$ is even, we first apply conjugation by $\check{\rho}(u)\in G_{\mathrm{ad}}(\!(u)\!)$ to $(p_{-1}+\sum_{i=1}^{n}u^{-d_{i}-m\lfloor\frac{d_{i}N}{m}\rfloor}c_{i}(u^{m})p_{i})\mathrm{d}u$, then apply $h(\frac{N}{2})$, we arrive at (61) as well.
The leading term of (61) is
(62) $Z=m(f+\sum_{m|d_{i}N}c_{i}(0)p_{i}^{\prime})\in\mathfrak{g}_{-N}.$
It suffices to show that for generic $c_{i}(0)$, the form (61) can be conjugated into $\mathfrak{j}^{+,\perp}$. Since $\mathfrak{g}_{-N}$ contains regular semisimple element $Y$, its regular semisimple locus is open and nonempty, so that we can assume $Z$ is regular semisimple. Observe that $\mathfrak{t}_{Y,-N}$ is a Cartan subspace of $\mathfrak{g}_{-N}$. Thus $Z$ can be conjugated into $\mathfrak{t}_{Y,-N}$. Without loss of generality, we assume $Z\in\mathfrak{t}_{Y,-N}$. Applying conjugation by $\exp(\tau_{k}u^{k})$ for $k\geq 1$ in the same way as in the first step of the proof of Lemma 20, we conjugate (61) into $\prod_{i\geq-N}\mathfrak{t}_{Y,i}u^{i}\frac{\mathrm{d}u}{u}\subset\mathfrak{j}^{+,\perp}$. This completes the proof. ∎
Denote
(63) $\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}:=\mathrm{Spec}\mathbb{C}[h_{i,j}\mid-N\leq\ell_{i,j}\leq-1].$
Note that $\mathrm{gr}A\simeq\mathrm{Fun}~{}\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$. We have projection $\mathrm{Hit}(D)_{\mathfrak{j}^{+}}\twoheadrightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$ using basis $h_{i,j}$.
Proposition 31.
  • (i)
    The composition $\mathfrak{j}^{+,\perp}\rightarrow\mathrm{Hit}(D)_{\mathfrak{j}^{+}}\rightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}^{*}}$ factors through quotient
    $\mathfrak{j}^{+,\perp}\rightarrow\mathfrak{j}^{+,\perp}/\mathfrak{j}^{\perp}\simeq\bar{\mathfrak{j}}^{*}.$
    The associated graded of the commutative diagram (55) is
    (64) ${\mathfrak{j}^{+,\perp}}$${\bar{\mathfrak{j}}^{*}}$${\mathrm{Hit}(D)_{\mathfrak{j}^{+}}}$${\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}}$
  • (ii)
    Let $\bar{\mathfrak{j}}^{*,\mathrm{rs}}\subset\bar{\mathfrak{j}}^{*}$ be the open locus consisting of elements whose components in $\mathfrak{t}_{Y,N}^{*}\simeq\mathfrak{t}_{Y,-N}$ are regular semisimple. Let $\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}\subset\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$ be the open locus whose coefficients $h_{i,j}$ satisfy that the vector (62) $Z=m(f+\sum_{\ell_{i,j}=-N}h_{i,j}p_{i}^{\prime})$ is regular semisimple. Then $h^{cl}(\bar{\mathfrak{j}}^{*,\mathrm{rs}})=\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$.
  • (iii)
    The restriction $h^{cl}:\bar{\mathfrak{j}}^{*,\mathrm{rs}}\rightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$ is étale with finite fibers that are isomorphic to the free conjugacy classes of the regular semisimple coefficient in $\mathfrak{t}_{Y,-N}$ under the finite little Weyl group $W_{0}:=N_{G_{0}}(\mathfrak{t}_{Y,-N})/C_{G_{0}}(\mathfrak{t}_{Y,-N})$.
Proof.
(i): The forms in $\mathfrak{j}^{+,\perp}$ whose leading terms in $\mathfrak{t}_{Y,-N}u^{-N}\frac{\mathrm{d}u}{u}$ are regular semisimple make up an open subspace of $\mathfrak{j}^{+,\perp}$. For such a form $\omega$, it can be conjugated into Kostant section (61), where the leading $N$ coefficients $\prod_{-N\leq i\leq-1}\mathfrak{g}_{i}u^{i}\frac{\mathrm{d}u}{u}$ of (61) are exactly given by the projection of $\omega$ to $\bar{\mathfrak{j}}\simeq\bar{\mathfrak{j}}^{*}$. This concludes the desired factorization and the commutative diagram.
(ii): This is clear from the discussion in part (i).
(iii): Let $\phi=(h_{i,j})_{-N\leq\ell_{i,j}\leq-1}\in\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$. As in part (i), the fiber $h^{cl,-1}(\phi)$ consists of elements $\tilde{\phi}=\sum_{i=1}^{N}\tilde{\phi}_{i}u^{i}\in\bar{\mathfrak{j}}$ and are conjugated to the Kostant section (61) of $\phi$. By the same proof as Proposition 23.(iii), elements in $h^{cl,-1}(\phi)$ are determined by their components in $\mathfrak{t}_{Y,-N}$. Thus $h^{cl,-1}(\phi)$ is isomorphic to the elements in $\mathfrak{t}_{Y,-N}$ that are conjugated to the regular semisimple leading term $Z$ in part (ii). By [Panyushev, Theorem 4.6.(i)] and [Vinberg, Theorem 2], $h^{cl,-1}(\phi)$ is isomorphic to $W_{0}\cdot Z$. Since the action of $W_{0}$ on $\mathfrak{t}_{Y,N}$ is given by Weyl group [Panyushev, Theorem 4.6.(ii)], it has free action on regular semisimple elements. Note that $W_{0}$ is finite by [Vinberg, Proposition 8].
We conclude that $h^{cl}:\bar{\mathfrak{j}}^{*,\mathrm{rs}}\rightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$ is isomorphic via Kostant section with a free quotient by $W_{0}$. ∎
Corollary 32.
  • (i)
    The maps $A^{\prime}\rightarrow\mathfrak{Z}_{\mathfrak{j}^{+}}$ and $A^{\prime}\rightarrow U(\bar{\mathfrak{j}})$ are injections.
  • (ii)
    The algebra $A^{\prime}$ is isomorphic to the intersection of $U(\bar{\mathfrak{j}})$ and $\mathfrak{Z}_{\mathfrak{j}^{+}}$ in $\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}}$, i.e. $A^{\prime}\simeq A$ in (33).
  • (iii)
    There is an isomorphism $A\simeq\mathrm{Fun}~{}\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$ that identifies $\mathrm{Spec}U(\bar{\mathfrak{j}})\rightarrow\mathrm{Spec}A$ with $\bar{\mathfrak{j}}\rightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$.
  • (iv)
    Let $(\mathrm{Spec}A)^{\mathrm{rs}}\subset\mathrm{Spec}A$ be the open subscheme of elements whose coefficients $S_{i,j}$ correspond under Feigin-Frenkel isomorphism to those coefficients $v_{i,j}$ such that $p_{-1}+\sum_{\ell_{i,j}=-N}v_{i,j}p_{i}$ is regular semisimple. Then $(\mathrm{Spec}A)^{\mathrm{rs}}\simeq\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$ under the isomorphism in (iii).
Proof.
(i): The injectivity of $A^{\prime}\rightarrow\mathfrak{Z}_{\mathfrak{j}^{+}}$ and $A^{\prime}\rightarrow U(\bar{\mathfrak{j}})$ follows from the surjectivity of $\mathrm{Hit}(D)_{\mathfrak{j}^{+}}\rightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$ and $\bar{\mathfrak{j}}\rightarrow\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$ of the $\mathrm{Spec}$ of associated graded algebras.
(ii): It suffices to prove for associated graded algebras, i.e. any regular function on $\mathfrak{j}^{+,\perp}$ that factors through both $\mathrm{Hit}(D)_{\mathfrak{j}^{+}}$ and $\bar{\mathfrak{j}}$ must factor through $\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$. Moreover, it suffices to prove over the open locus of $\mathfrak{j}_{0}^{+,\perp}$ with regular semisimple leading term.
Indeed, let $f$ be a regular function on $\mathfrak{j}_{0}^{+,\perp}$ that factors through the local Hitchin map and is constant over cosets of $\mathfrak{j}^{\perp}$. The local Hitchin map on $\mathfrak{j}_{0}^{+,\perp}$ can be computed by first conjugate the form into $\prod_{i\geq-N}\mathfrak{t}_{Y,i}u^{i}\frac{\mathrm{d}u}{u}$, next conjugate it into Kostant section (61), then just take the coefficients of Kostant section to be the values of $h_{i,j}$. By $\mathfrak{j}^{\perp}$-invariance, the values of $f$ depend on the first $N$ leading terms of the Kostant section. Thus $f$ factors through $\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}$.
(iii): We just need to show that $A$ is isomorphic to its associated graded algebra. In fact, we have seen in the proof of Proposition 29 that the action of $S_{i,j}$ on $1\in\mathrm{Vac}_{\mathfrak{j}^{+}}$ factors through the projection to the length $d_{i}$ tensors in $S_{i,j}$. Since the action gives an embedding of $A$ by (i), we get $A\simeq\mathrm{gr}A$ as algebras.
(iv): Recall the Kostant sections in $\check{\mathfrak{g}}$ and $\mathfrak{g}$ are identified via $\check{\mathfrak{g}}/\!\!/\check{G}\simeq\mathfrak{g}^{*}/\!\!/G\simeq\mathfrak{g}/\!\!/G$. Also note that the tuples $(f,p_{i}^{\prime})$ and $(p_{-1},p_{i})$ are conjugated. The statement follows from Proposition 31.(ii). ∎
6.5. Central support of $\mathrm{Vac}_{\mathfrak{j}^{+}}$
Let $\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\leq\nu}$ be the space of opers on $D^{\times}$ with slope bounded above by $\nu=\frac{N}{m}$. By the oper slope formula (14),
$\mathrm{Fun}~{}\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\leq\nu}=\mathbb{C}[S_{i,j}\mid\ell_{i,j}\geq-N].$
Denote $\mathrm{Op}(D)_{\mathfrak{j}^{+}}=\mathrm{Spec}\mathfrak{Z}_{\mathfrak{j}^{+}}\hookrightarrow\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})$.
Lemma 33.
$\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\mathfrak{j}^{+}}=\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\leq\nu}$.
Proof.
For any $S_{i,j}$ with $\ell_{i,j}<-N$, write $S_{i,j}$ as linear combination of tensors as in the proof of Proposition 29. We have
$\sum_{r=1}^{k}a_{r}=m(j-d_{i}+1)=(d_{i}-1)N-\ell_{i,j}>d_{i}N.$
Since $k\leq d_{i}$, $a_{1}\leq\cdots\leq a_{k}$, we have $a_{k}\geq\frac{\sum_{r=1}^{k}a_{r}}{d_{i}}>N$. Therefore $S_{i,j}\cdot 1=0$ in $\mathrm{Vac}_{\mathfrak{j}^{+}}$. The action of $\mathfrak{Z}$ on $\mathrm{Vac}_{\mathfrak{j}^{+}}$ factors as
$\mathfrak{Z}\rightarrow\mathrm{Fun}~{}\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\leq\nu}\twoheadrightarrow\mathfrak{Z}_{\mathfrak{j}^{+}}\hookrightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j}^{+}}.$
By Proposition 30, the associated graded of the middle surjection in the above is an isomorphism:
$\mathrm{gr}\mathrm{Fun}~{}\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\leq\nu}=\mathrm{Fun}~{}\mathrm{Hit}(D)_{\mathfrak{j}^{+}}^{\prime}=\mathrm{Fun}~{}\mathrm{Hit}(D)_{\mathfrak{j}^{+}}=\mathrm{gr}\mathfrak{Z}_{\mathfrak{j}^{+}}.$
The lemma follows. ∎
7. Isoclinic local geometric Langlands
We construct a correspondence between irreducible isoclinic formal connections and toral supercuspidal representations in the sense of Frenkel-Gaitsgory [FGLocal].
7.1. Matching canonical forms with $K$-types
We first match the irreducible isoclinic formal connections with pairs $(J,\tilde{\phi})$.
7.1.1.
Firstly, the matching of the leading terms of the isoclinic formal connection and $\tilde{\phi}$ are given by [CYTheta, Lemma 3]. Explicitly, recall $X\in\check{\mathfrak{g}}_{-N}$ and $Y\in\mathfrak{g}_{N}^{*}$ are regular semisimple elements. Denote by $\check{G}_{0}$(resp. $G_{0}$) the subgroup of $\check{G}$(resp. $G$) with Lie algebra $\check{\mathfrak{g}}_{0}$(resp. $\mathfrak{g}_{0}$).
Lemma 34.
There exists an isomorphism
(65) $\mathfrak{g}_{N}^{*}/\!\!/G_{0}\simeq\check{\mathfrak{g}}_{-N}/\!\!/\check{G}_{0}$
that restricts to a bijection between regular semisimple conjugacy classes
(66) $\mathfrak{g}_{N}^{*,\mathrm{rs}}/G_{0}\simeq\check{\mathfrak{g}}_{-N}^{\mathrm{rs}}/\check{G}_{0}.$
Proof.
Note that the longest element of Weyl group $w_{0}\in W$ conjugates $\check{\rho}$ to $-\check{\rho}$. Thus $\theta=\mathrm{Ad}_{\check{\rho}(\zeta_{m})}$ and $\theta^{-1}$ are conjugated by $\mathrm{Ad}_{w_{0}}$. This induces an isomorphism $\check{\mathfrak{g}}_{-N}\simeq\check{\mathfrak{g}}_{N}$. The isomorphism follows from replacing $\zeta_{m}$ with $\zeta_{m}^{N}$ in [CYTheta, Lemma 3], where regular semisimple eigenvectors are the same as stable vectors by Lemma 9 of loc. cit.. Here although the grading on $\mathfrak{g}$ is defined by $\Theta=\mathrm{Ad}_{\lambda(\zeta_{m})}$ rather than $\mathrm{Ad}_{\check{\rho}(\zeta_{m})}$ in loc. cit., since $\lambda$ is conjugated to $\check{\rho}$, the $\zeta_{m}^{N}$-eigenspaces of the gradings by $\Theta$ and $\mathrm{Ad}_{\check{\rho}(\zeta_{m})}$ are conjugated. ∎
7.1.2.
Fix $X\in\check{\mathfrak{g}}_{-N}^{\mathrm{rs}}$ and pick any irreducible isoclinic formal connection $\nabla\in\mathrm{Loc}(X,\nu)$. Let $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)$ be any oper above $\nabla$. By Proposition 23.(iii), the values of $\chi$ on $S_{i,j}$ for $(i,j)\in\tilde{A}_{\nu}$ are uniquely determined by $\nabla$, independent of the choice of $\chi$. This determines a character $\phi$ of $A$ that lands inside $(\mathrm{Spec}A)^{\mathrm{rs}}$. By (iii) and (iv) of Corollary 32, we can regard $\phi$ as a point of $\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$. By Proposition 31.(ii), $\phi$ can be extended to a character $\tilde{\phi}$ of $\bar{\mathfrak{j}}$ that corresponds to a point in $\bar{\mathfrak{j}}^{*,\mathrm{rs}}\simeq\bar{\mathfrak{j}}^{\mathrm{rs}}$, i.e. $\tilde{\phi}\in h^{cl,-1}(\phi)$. Inflated by $J\rightarrow J/J^{+}\simeq\bar{\mathfrak{j}}$ and composed with exponential, we obtain a character of $J$ of depth $\nu$, which we still denote by $\tilde{\phi}$.
Denote $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{j}+\mathbb{C}\mathbf{1}}\tilde{\phi}$ and $\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}=\mathrm{Im}(\mathfrak{Z}\rightarrow\mathrm{End}\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}})$. Let $\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi}):=\mathrm{Spec}\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}}$. Denote by $\mathrm{Op}_{\check{\mathfrak{g}}}(\nabla)=p^{-1}(\nabla)$ the opers with underlying connection $\nabla$.
Lemma 35.
  • (i)
    Let $\psi$ be the character on $A_{-N}:=\mathbb{C}[S_{i,j},\ell_{i,j}=-N]\subset A^{\prime}\simeq A$ determined by the coefficients of the unique element $p_{-1}+\sum_{\ell_{i,j}=-N}v_{i,j}p_{i}\in\check{G}\cdot X$. Then $\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)$ is the central support of $\mathrm{Vac}_{\mathfrak{j}^{+}}/\ker\psi$.
  • (ii)
    For any $\tilde{\phi}_{1},\tilde{\phi}_{2}\in h^{cl,-1}(\phi)$, $\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi}_{1})=\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi}_{2})$.
  • (iii)
    $\mathrm{Op}_{\check{\mathfrak{g}}}(\nabla)=\bigcup_{\tilde{\phi}\in h^{cl,-1}(\phi)}\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})=\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})$ for any $\tilde{\phi}\in h^{cl,-1}(\phi)$.
Proof.
(i): This follows immediately from Lemma 33 and Proposition 23.(i).
(ii): We first exhibit the central support in a more explicit form. For $\tilde{\phi}=\tilde{\phi}_{1}$ or $\tilde{\phi}_{2}$,
$\mathrm{End}\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}\simeq\mathrm{Hom}_{\mathfrak{j}}(\mathbb{C}_{\tilde{\phi}},\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}})\simeq\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}^{\ker\tilde{\phi}}\subset\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}=\widehat{U}(\mathfrak{g})/\widehat{U}(\mathfrak{g})\ker\tilde{\phi}.$
Thus
$\ker(\mathfrak{Z}\rightarrow\mathfrak{Z}_{\mathfrak{j},\tilde{\phi}})=\ker(\mathfrak{Z}\rightarrow\widehat{U}(\mathfrak{g})/\widehat{U}(\mathfrak{g})\ker\tilde{\phi})=\mathfrak{Z}\cap(\widehat{U}(\mathfrak{g})\ker\tilde{\phi}).$
By Proposition 31.(iii) and its proof, any two elements $\tilde{\phi}_{1},\tilde{\phi}_{2}\in h^{cl,-1}(\phi)$ satisfy $\tilde{\phi}_{2}=\mathrm{Ad}_{w}\tilde{\phi}_{1}$ for some element $w\in N_{G_{0}}(\mathfrak{t}_{Y,-N})$, so that $\ker\tilde{\phi}_{2}=\mathrm{Ad}_{w}\ker\tilde{\phi}_{1}\subset U(\mathfrak{j})$. Since $\mathrm{Ad}_{w}$ acts trivially on $\mathfrak{Z}$, we obtain
$\mathfrak{Z}\cap(\widehat{U}(\mathfrak{g})\ker\tilde{\phi}_{1})=\mathrm{Ad}_{w}(\mathfrak{Z}\cap(\widehat{U}(\mathfrak{g})\ker\tilde{\phi}_{1}))=\mathfrak{Z}\cap(\widehat{U}(\mathfrak{g})\ker\tilde{\phi}_{2}).$
From the above discussion, $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}_{1}}$ and $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}_{2}}$ have the same central supports.
(iii): The second equality follows from (ii). To prove the first equality, we first show inclusion $\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})\subset\mathrm{Op}_{\check{\mathfrak{g}}}(\nabla)$ for any $\tilde{\phi}\in h^{cl,-1}(\phi)$. Let $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})$. Note that we have surjection
$q:\mathrm{Vac}_{\mathfrak{j}^{+}}=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{j}+\mathbb{C}\mathbf{1}}U(\bar{\mathfrak{j}})\rightarrow\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{j}+\mathbb{C}\mathbf{1}}\tilde{\phi}.$
Thus $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\mathfrak{j}^{+}}=\mathrm{Op}_{\check{\mathfrak{g}}}(D)_{\leq\nu}$. Also, by Proposition 29 $A$ acts on $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$ via $\tilde{\phi}$. By Proposition 23.(iii), the underlying connection of $\chi$ is $\nabla$.
Conversely, let $\phi$ be the character of $A$ that is determined by $\nabla$ via Proposition 23.(iii). By Lemma 33 and Proposition 29, $\mathrm{Op}_{\check{\mathfrak{g}}}(\nabla)$ is the central support of
$\mathrm{Vac}_{\phi}:=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{j}+\mathbb{C}\mathbf{1}}(U(\bar{\mathfrak{j}})/\ker\phi).$
Regard $\phi\in\mathrm{Hit}(D)_{\bar{\mathfrak{j}}}^{\mathrm{rs}}$. Then $U(\bar{\mathfrak{j}})/\ker\phi\simeq\mathrm{Fun}~{}h^{cl,-1}(\phi)$. By Proposition 31.(iii), $|h^{cl,-1}(\phi)|=|W_{0}|$ is finite. We obtain
$U(\bar{\mathfrak{j}})/\ker\phi\simeq\bigoplus_{\tilde{\phi}\in h^{cl,-1}(\phi)}U(\bar{\mathfrak{j}})/\ker\tilde{\phi}$
and
$\mathrm{Vac}_{\phi}\simeq\bigoplus_{\tilde{\phi}\in h^{cl,-1}(\phi)}\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\bar{\mathfrak{j}}+\mathbb{C}\mathbf{1}}\mathbb{C}_{\tilde{\phi}}.$
Thus the central support of $\mathrm{Vac}_{\phi}$ is the union of central supports of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$’s. ∎
7.2. The $K$-type of the central block
For an oper $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(D^{\times})$, denote by $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ the category of smooth $\widehat{\mathfrak{g}}$-modules on which the center $\mathfrak{Z}$ acts via $\chi$. We say a $\widehat{\mathfrak{g}}$-module is $(J,\tilde{\phi})$-equivariant if it contains an object mapped into by a nonzero map from $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$. Since $J$ is pro-unipotent, this is equivalent to that it contains a nonzero $\ker\tilde{\phi}$-equivariant object.
We have the following refinement of Corollary 15 in the case of isoclinic formal connections.
Theorem 36.
Let $Y\in\mathfrak{g}_{N}^{*,\mathrm{rs}}$ and $X\in\check{\mathfrak{g}}_{-N}^{\mathrm{rs}}$ match under (66). The category $\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ contains a nonzero $(J,\tilde{\phi})$-equivariant object for some $\tilde{\phi}\in\bar{\mathfrak{j}}^{*,\mathrm{rs}}$ with component $Y$ in $\mathfrak{t}_{Y,N}$ if and only if $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)$, i.e. the underlying connection $\nabla$ of $\chi$ is isoclinic of depth $\nu$ with leading term $X$.
Moreover, $\tilde{\phi}$ can be any preimage of $\phi\in\mathrm{Spec}A$ for the $\phi$ associated to $\nabla$. Conversely, $\nabla$ has a unique equation of the form (43) where $v_{i,j}=\tilde{\phi}(\overline{S_{i,j}})$ for $\overline{S_{i,j}}$ the image of $S_{i,j}$ in $A\subset U(\bar{\mathfrak{j}})$, $-N\leq\ell_{i,j}\leq-1$.
Proof.
If part: Let $\nabla$ be the underlying connection of $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu)$. Let $\phi\in\mathrm{Spec}A$ correspond to $\nabla$. By Lemma 35, $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})$ for every $\tilde{\phi}\in h^{cl,-1}(\phi)$. As in Corollary 15,
$\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}/\ker\chi$
is nonzero $(J,\tilde{\phi})$-equivariant.
Only if part: by the same argument as the proof of [FrenkelBook, Lemma 10.3.2], there exists a module $M\in\widehat{\mathfrak{g}}-\mathrm{mod}_{\chi}$ with a nonzero map $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}\rightarrow M$. Then
$\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{j},\tilde{\phi})\subset\mathrm{Op}_{\check{\mathfrak{g}}}(\nabla)\subset\mathrm{Op}_{\check{\mathfrak{g}}}(X,\nu).$
The description of $\tilde{\phi}$ and $\nabla$ in terms of each other is clear from Lemma 35 and its proof. ∎
Remark 37.
To sum up, for toral supercuspidal representations and irreducible isoclinic connections, Conjecture.(i)-(iii) 14 has been established in Lemma 35 and Theorem 36. The part (iv) of the conjecture follows from the same proof as Lemma 35.(ii).
8. Application: Airy connections
As an example and application of our local results, we prove that the Langlands parameters of Hecke eigensheaves constructed in [JKY] are Airy connections for reductive groups [KSRigid, HJRigid].
8.1. Airy connections and Airy automorphic data
8.1.1. Airy connections
Following [HJRigid, Definition 5.5.1], an Airy $\check{G}$-connection is a $\check{G}$-connection on $\mathbb{P}^{1}-\{0\}$ [^4] We let the singularity be at 00 instead of $\infty$ for convenience of notations. that is isoclinic of slope $\nu=\frac{1+h}{h}$ at 00. Here $h$ is the Coxeter number of $\check{G}$.
For an Airy connection $\nabla$, its restriction $\nabla_{0}:=\nabla|_{D_{0}^{\times}}$ is an isoclinic formal connection. Moreover, since $h$ is a regular elliptic number, by the same proof as [YiFG, Lemma 7] we know $\nabla_{0}$ is irreducible. By [JYDeligneSimpson, Theorem 1.8, Theorem 4.1], any isoclinic formal connection $\nabla_{0}$ of slope $\nu$ is the formal type at 00 of an Airy connection $\nabla$. By [HJRigid, Theorem 5.5.2], Airy connections are physically rigid, i.e. for any Airy connection $\nabla$ and another $\check{G}$-connection $\nabla^{\prime}$ on $\mathbb{P}^{1}-\{0\}$, $\nabla|_{D_{0}^{\times}}\simeq\nabla^{\prime}|_{D_{0}^{\times}}$ implies $\nabla\simeq\nabla^{\prime}$. Therefore Airy connections are in bijection with isoclinic formal connections of slope $\nu$. Denote by $\mathrm{Ai}_{\check{G}}(X)$ the isomorphism classes of Airy connections with leading term $X$ in its canonical form at 00.
Since $\nu>1$, combining the above with (43) and Proposition 25, we obtain
Proposition 38.
We have isomorphisms
$\mathrm{Ai}_{\check{G}}(X)\simeq\mathrm{Loc}(X,\frac{1+h}{h},\mathbb{P}^{1}-\{0\})\simeq\mathrm{Loc}(X,\frac{1+h}{h}),$
where each such Airy connection has an unique equation of the form
$\mathrm{d}+(t^{-2}p_{-1}+v_{n,2h}t^{-3}p_{n}+\sum_{i=1}^{n}v_{i,2d_{i}-1}t^{-2}p_{i})\mathrm{d}t,\quad p_{-1}+v_{n,2h}p_{n}\in\check{G}\cdot X.$
Comparing the above with [KSMon, equation (8)], we conclude that every Airy $\check{G}$-connection is framable. Observe that (7) is a special case of the above equation.
8.1.2. Airy automorphic data
The Airy automorphic data [JKY, Definition 8] are special cases of toral $K$-types $(J,\tilde{\phi})$ of depth $\nu=\frac{1+h}{h}$ defined in §6.1555The use of notations $\tilde{\phi},\phi$ are slightly different from [JKY].. Here $m=h$, $N=1+h$. We recall the particular choice of $\tilde{\phi}$ in the loc. cit., which we will see being necessary. Recall $\tilde{\phi}\in\bar{\mathfrak{j}}^{*}\simeq\bigoplus_{-1-h\leq i\leq-1}\mathfrak{t}_{Y,i}u^{i}\frac{\mathrm{d}u}{u}$, where $Y\in\mathfrak{g}_{-1-h}^{\mathrm{rs}}$ is the component of $\tilde{\phi}$ in $\mathfrak{g}_{-1-h}$. We require $\tilde{\phi}$ to have zero components in $\mathfrak{t}_{Y,i}$ for $-h\leq i\leq-(1+\frac{h}{2})$ when $h$ is even (resp. for $-h\leq i\leq-(n+1)$ when $h=2n+1$ is odd, $\mathfrak{g}=\mathfrak{sl}_{2n+1}$). We call such $\tilde{\phi}$ special.
Denote by $\mathcal{G}^{\prime}$ the group scheme on $\mathbb{P}^{1}$ satisfying $\mathcal{G}^{\prime}|_{\mathbb{P}^{1}-\{0\}}\simeq G\times(\mathbb{P}^{1}-\{0\})$ and $\mathcal{G}(\mathcal{O}_{0})=J^{+}$. Denote by $\mathrm{Bun}_{\mathcal{G}^{\prime}}$ the moduli stack of $\mathcal{G}^{\prime}$-bundles. The group $J$ acts on $\mathrm{Bun}_{\mathcal{G}^{\prime}}$ at 00. Let $\mathcal{L}_{\tilde{\phi}}=\tilde{\phi}^{*}(\mathrm{d}-\mathrm{d}t)$ be a character $D$-module on $J$.
Proposition 39.
[JKY, Lemma 21] For $\tilde{\phi}$ special, there exists unique $(J,\mathcal{L}_{\tilde{\phi}})$-equivariant irreducible holonomic $D$-module $\mathcal{A}_{\tilde{\phi}}$ on $\mathrm{Bun}_{\mathcal{G}^{\prime}}$ up to isomorphism.
We explain below why the assumption of $\tilde{\phi}$ being special is necessary in Proposition 39.
Recall that relevant orbits are $J$-orbits on $\mathrm{Bun}_{\mathcal{G}^{\prime}}$ over which $(J,\mathcal{L}_{\tilde{\phi}})$-equivariant sheaves can support. In other words, these are orbits over which $\mathcal{L}_{\tilde{\phi}}$ is trivial on the stabilizer of $J$. The proof in [JKY, §4.2.1,§4.2.2] actually goes through without assuming $\tilde{\phi}$ being special, which shows that there is at most one relevant orbit (assuming $G$ is simply-connected): the orbit of trivial $\mathcal{G}^{\prime}$-bundle. However, we have the following observation:
Proposition 40.
The unit orbit is relevant if and only if $\tilde{\phi}$ is special.
Proof.
We assume $h$ even. The proof for $h$ odd is just replacing all the $\frac{h}{2}$ with $n+1$, $h=2n+1$.
Denote by $\mathfrak{g}=\bigoplus_{i\in\mathbb{Z}}\mathfrak{g}(i)$ the $\mathbb{Z}$-grading defined by $\check{\rho}_{G}$. Then $\mathfrak{g}_{i}=\mathfrak{g}(i)\oplus\mathfrak{g}(i+h)$ for $-h-1\leq i\leq-1$. Regard $\tilde{\phi}\in\bar{\mathfrak{j}}^{*}\simeq\bigoplus_{-1-h\leq i\leq-1}\mathfrak{t}_{Y,i}u^{i}\frac{\mathrm{d}u}{u}$. By [JKY, §4.4], for the unit orbit to be relevant, we need
(67) $\mathrm{Res}(\kappa(\tilde{\phi},\bigoplus_{i=1+\frac{h}{2}}^{h-1}\mathfrak{g}(i)))\frac{\mathrm{d}t}{t}\simeq\mathrm{Res}(\kappa(\tilde{\phi},\bigoplus_{i=1+\frac{h}{2}}^{h-1}\mathfrak{g}(i)u^{i}))\frac{\mathrm{d}u}{u}=0.$
Let $\tilde{\phi}_{i}=Y_{i}u^{i}\frac{\mathrm{d}u}{u}$ be the component of $\tilde{\phi}$ in $\mathfrak{t}_{Y,i}u^{i}\frac{\mathrm{d}u}{u}$. Then (67) holds if and only if $\kappa(Y_{-i},\mathfrak{g}(i))=0$ for $11+\frac{h}{2}\leq-i\leq h-$ 1 (note $Y_{-h}\in\mathfrak{t}_{Y,-h}=\mathfrak{t}_{Y,0}=0$). Denote by $p_{-}$ the projection from $\mathfrak{t}_{Y,-i}\subset\mathfrak{g}_{-i}$ to $\mathfrak{g}(-i)$. Then
$\kappa(Y_{-i},\mathfrak{g}(i))=\kappa(p_{-}(Y_{-i}),\mathfrak{g}(i))=\kappa(p_{-}(Y_{-i}),\mathfrak{g}_{i}).$
By [JKY, Lemma 15], $p_{-}$ is injective. Since $\kappa$ is non-degenerate on $\mathfrak{g}_{-i}\times\mathfrak{g}_{i}$, $\kappa(p_{-}(Y_{-i}),\mathfrak{g}_{i})=0$ if and only if $p_{-}(Y_{-i})=0$ if and only if $Y_{-i}=0$. ∎
8.2. The Langlands correspondence
Theorem 41.
Let $(J,\tilde{\phi})$ be such that $\tilde{\phi}$ is special. Let $\nabla_{0}$ be the isoclinic formal connection corresponding to $\tilde{\phi}$ as in Theorem 36. Let $\nabla$ be the Airy connection with formal type $\nabla_{0}$ at 00.
  • (i)
    $\nabla$ is the eigenvalue of $\mathcal{A}_{\tilde{\phi}}$.
  • (ii)
    $\nabla$ is physically rigid.
Proof.
The proof is just a simplified version of the proof of [CYTheta, Theorem 5]. We sketch it for completeness. We will prove (i) and (ii) together.
Let $\nabla^{\prime}$ be a $G$-connection on $\mathbb{P}^{1}-\{0\}$ such that $\nabla^{\prime}|_{D_{0}^{\times}}\simeq\nabla_{0}$. Recall $\nabla_{0}$ is irreducible, so is $\nabla^{\prime}$. By [Arinkin, Corollary 1.1], $\nabla^{\prime}$ has an oper structure $\chi\in\mathrm{Op}_{\check{\mathfrak{g}}}(U)$ over some open curve $U=\mathbb{P}^{1}-S$ where $S=\{z_{0}=0,z_{1},...,z_{q}\}$. Here if $q=0$, we assign $z_{1}=\infty$. So we always assume $q>0$.
At $z=z_{i}$, $i>0$, $\chi_{i}:=\chi|_{D_{z_{i}}^{\times}}$ is a monodromy free oper. As in [YiFG, §4.2.2], $\chi_{i}$ lands inside the central support $\mathrm{Op}_{\check{\mathfrak{g}}}^{\mathrm{reg}}(D_{z_{i}})_{\varpi(-\lambda_{i}-\rho)}$ of Weyl module $\mathbb{V}_{\lambda_{i}}=\mathrm{Ind}^{\widehat{\mathfrak{g}}}_{\mathfrak{g}[\![t]\!]+\mathbb{C}\mathbf{1}}V_{\lambda_{i}}$, where $\lambda$ is a dominant integral weight of $\mathfrak{g}$, $V_{\lambda_{i}}$ is the irreducible highest weight $\lambda_{i}$ representation of $\mathfrak{g}$.
At $z_{0}=0$, by Lemma 35.(iii), $\chi_{0}=\chi|_{D_{0}^{\times}}$ lands inside the central support $\mathrm{Op}_{\check{\mathfrak{g}}}(\mathfrak{h},\tilde{\phi})$ of $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}$ for every $\tilde{\phi}\in h^{cl,-1}(\phi)$ where $\phi\in\mathrm{Spec}A$ corresponds to $\nabla_{0}$ under (43).
Let $\mathrm{Loc}$ be the localization functor as in [YiFG, §3]. Let $\omega_{\mathrm{Bun}_{\mathcal{G}^{\prime}}}^{-1/2}$ be a square root line bundle of the canonical sheaf of $\mathrm{Bun}_{\mathcal{G}^{\prime}}$ which exists by [Zhu, Proposition 7]. Consider
(68) $\mathcal{A}_{\chi}:=\omega_{\mathrm{Bun}_{\mathcal{G}^{\prime}}}^{-1/2}\otimes\mathrm{Loc}(\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}\otimes_{\mathbb{C},i}\mathbb{V}_{\lambda_{i}}/\ker\chi_{i}).$
Note $\mathrm{Vac}_{\mathfrak{j},\tilde{\phi}}=\mathrm{Vac}_{\mathfrak{j}^{+}}/\ker\tilde{\phi}$. By construction, $\mathcal{A}_{\chi}$ is $(J,\tilde{\phi})$-equivariant that is Hecke eigen with eigenvalue $\nabla^{\prime}$. It only remains to show that $\mathcal{A}_{\chi}$ is nonzero and holonomic, so that it is a nonzero complex of copies of the unique irreducible $(J,\tilde{\phi})$-equivariant holonomic $D$-module in Proposition 39.
Let $L^{-}G$ be the opposite loop group, then $\mathrm{Bun}_{\mathcal{G}^{\prime}}\simeq[L^{-}G\backslash LG/J^{+}]$. By [JKY, Theorem 13, Proposition 19], the unique relevant orbit $j:O=[L^{-}G\backslash L^{-}GJ/J^{+}]\hookrightarrow\mathrm{Bun}_{\mathcal{G}^{\prime}}$ is locally closed. As in §6.5 of [CYTheta], $\mathcal{A}_{\chi}\simeq j_{*}j^{*}\mathcal{A}_{\chi}$ is holonomic.
It remains to show $\mathcal{A}_{\chi}\neq 0$. By [JKY, Proposition 19], we have open embedding $j^{\prime}:O^{\prime}=[L^{-}G\backslash L^{-}GI(1)/J^{+}]\hookrightarrow\mathrm{Bun}_{\mathcal{G}^{\prime}}$. To see $\mathcal{A}_{\chi}\neq 0$, it suffices to show the associated graded module $\mathrm{gr}H^{0}(j^{\prime*}\mathcal{A}_{\chi})$ is nonzero. Although here $O^{\prime}$ is the quotient by an unipotent group on an affine space unlike being an affine space in [CYTheta, Lemma 17], since it suffices to show that at least one stalk is non-vanishing, we can still use the proof in loc. cit. to reduce the problem to the non -vanishing of the following module:
$\mathcal{O}_{T^{*}O^{\prime}}\otimes_{\mathrm{Fun}~{}\mathrm{Hit}_{\mathcal{G}^{\prime}}\otimes_{A}\mathrm{Fun}~{}\bar{\mathfrak{j}}^{*}}\mathbb{C}_{0}.$
Here $\mathrm{Hit}_{\mathcal{G}^{\prime}}:=\mathrm{Hit}(\mathbb{P}^{1}-\{0\})\times_{\mathrm{Hit}(D_{0}^{\times})}\mathrm{Hit}(D)_{\mathfrak{j}^{+}}$. By Proposition 30, $\mathrm{Hit}_{\mathcal{G}^{\prime}}\simeq\bigoplus_{i=1}^{n}\Gamma(\mathcal{O}(\lfloor\frac{d_{i}}{h}\rfloor))\simeq\mathbb{A}^{n+1}$ is an affine space. The character $\mathbb{C}_{0}$ corresponds to the unique $\mathbb{G}_{m}$-fixed point of $\mathrm{Hit}_{\mathcal{G}^{\prime}}\times_{\mathrm{Spec}(A)}\bar{\mathfrak{j}}^{*}$.
It remains to show that the image of composition
$\mu:T^{*}O^{\prime}\hookrightarrow T^{*}\mathrm{Bun}_{\mathcal{G}^{\prime}}\rightarrow\bar{\mathfrak{j}}^{*}$
contains 00. Note $O^{\prime}=[U\backslash I(1)/J^{+}]$ where $U=L^{-}G\cap I(1)$. Let $\mu_{U}:T^{*}(I(1)/J^{+})\rightarrow\mathfrak{u}^{*}=\mathrm{Lie}(U)^{*}$ be the moment map. Then $T^{*}O^{\prime}\simeq U\backslash\mu_{U}^{-1}(0)$. Note that $\mu_{U}$ factors through the cotangent map $(I(1)/J^{+})^{*}\rightarrow\mathfrak{u}^{*}$ of group morphism $U\rightarrow I(1)/J^{+}$. Thus the quotient of the zero section of $T^{*}(I(1)/J^{+})$ by $U$ is contained in $T^{*}O^{\prime}$ and maps to 00 via $(I(1)/J^{+})^{*}\rightarrow\bar{\mathfrak{j}}^{*}$. This completes the proof. ∎
Remark 42.
For $\check{G}=\mathrm{GL}_{n}$, part (i) of Theorem 41 has been proved in [JKY, Corollary 36]. The physical rigidity of Airy connections for general $\check{G}$ and irregular types has been proved in [HJRigid, Theorem 5.5.2], and here we recover it only for those Airy connections corresponding to special $\tilde{\phi}$. It is still unclear to us how to generalize the Airy automorphic datum to cover all Airy connections in the eigenvalues.
We confirm Conjecture 29 of [JKY].
Corollary 43.
Let $(J,\tilde{\phi})$ be an Airy automorphic datum where $\tilde{\phi}$ is special. Denote the eigenvalue $\check{G}$-connection of the associated Hecke eigensheaf by $\nabla_{\tilde{\phi}}$.
  • (i)
    The irregularity of $\nabla_{\tilde{\phi}}^{\mathrm{Ad}}$ at 00 is $n(h+1)$.
  • (ii)
    The local differential Galois group $\mathcal{I}_{0}\subset\check{G}$ of $\nabla_{\tilde{\phi}}$ at 00 satisfies $\check{\mathfrak{g}}^{\mathcal{I}_{0}}=0$.
  • (iii)
    $\nabla_{\tilde{\phi}}$ is cohomologically rigid.
Proof.
We know from Theorem 41 that $\nabla_{\tilde{\phi}}$ is an Airy connections. For Airy connections, (i)-(iii) are proved in [JYDeligneSimpson, Lemma 6.2, Proposition 6.4]. ∎